### v1

# SkyMask: Attack-agnostic Robust Federated Learning with Fine-grained Learnable Masks

Abstract
Federated Learning (FL) is becoming a popular paradigm for leveraging distributed data and preserving data privacy. However, due to the distributed characteristic, FL systems are vulnerable to Byzantine attacks that compromised clients attack the global model by uploading malicious model updates. Most existing Byzantine-robust FL systems statistically analyze the weights of whole individual model updates uploaded by clients to defend against Byzantine attacks. With the development of layer-level and parameter-level fine-grained attacks, the attacks’ stealthiness and effectiveness have been significantly improved. Due to unawareness or overreaction, the existing model-level defense methods degrade the training efficiency and model performance. To address this problem, we propose SkyMask, a new attack-agnostic robust FL system that leverages fine-grained learnable masks to identify malicious model updates at the parameter-level. Specifically, the FL server applies parameter-level masks to model updates uploaded by clients and trains the masks over a small clean dataset (i.e., root dataset) to learn the subtle difference between benign and malicious model updates in a high-dimension space. Our extensive experiments involve different models on three public datasets under state-of-the-art (SOTA) attacks, where the results show that SkyMask achieves up to 10% higher testing accuracy compared with SOTA defense strategies and successfully defends against attacks with malicious clients of a high fraction up to 80%. In the meantime, the experimental results demonstrate the scalability of our approach and the weak dependence on the data distribution of the root dataset.
I Introduction
Every user click is an oil drop of the digital era that powers up today’s information industry. Massive volumes of data generated by users and collected by online services (e.g., social networks, online sales, and streaming platforms) are used to train machine learning (ML) models (e.g., recommendation algorithms) that enhance user experience and boost profits. However, with increasingly strict data and privacy policies, such as the GDPR by EU and ATT by Apple, web services tracking and collecting user data for centralized training becomes inappropriate. The emergence of federated learning (FL) has addressed such conflicts between ML training and data privacy [1]. Instead of centralizing training data, FL distributes a model to clients, trains the model locally, and aggregates the local models into a global model without leaking any clients’ local data, which has been applied in many web applications, including online recommendation systems and personal healthcare.
However, due to FL’s distributed design, attackers can easily attack the FL system by compromising client participants and smuggling malicious model updates, known as Byzantine attacks [2, 3, 4, 5, 6, 7]. Based on whether the compromised clients poison their local data or models, Byzantine attacks can be classified into data poisoning and model poisoning attacks. Data poisoning attacks corrupt the global model by injecting contaminated training samples to the local dataset [2, 7]. Model poisoning attacks [3, 5] directly manipulate local model updates to compromise the global model. To tackle these attacks, researchers have explored various defense strategies, most of which leverage coarse-grained model-level statistics, distances, and cosine-similarities to detect outlier model updates (i.e., Multi-Krum [8] and FLTrust [9]) or greedily filter out outlier parameters (i.e., Trim [10] and Bulyan [11]).
Unfortunately, research on fine-grained attacking methodologies is becoming trendy due to their precise balancing between attacking stealthiness and effectiveness. Fine-grained model poisoning attacks [4, 12] leverage the varying sensitivities of different neural network layers and parameters to craft adaptive and stealthy attacks. FIBA [13] injects small pixel-wise perturbations in the frequency domain to conduct data poisoning. Existing model-level defense strategies either fail to prevent such fine-grained attacks or spill over the benign clients, which extensively sacrifices training efficiency and model quality.
This paper proposes SkyMask, a new Byzantine-robust FL system that defends against agnostic attacks with fine-grained learnable masks at the parameter-level. Specifically, after collecting local models from participant clients, the FL server freezes and multiplies each of the local model parameters with a learnable mask variable. Then, all masked local models are trained on a small and clean data set (i.e., the root dataset [9]) together. During the training on the root dataset, the mask variables applied to the frozen local model parameters learn to improve the model accuracy. Particularly, the masks applied to malicious models are trained to correct their potential misbehaviors at the parameter-level. Since utilizing fine-grained learnable masks does not enforce customization for specific attacks or refactoring the FL training objective functions, SkyMask gains three superior capabilities beyond existing defense strategies: 1) effective and efficient detection of fine-grained stealthy attacks, 2) attack-agnostic defense against a wide spectrum of attack methods, and 3) strong compatibility and complementarity with existing FL and defense methodology. Our main contributions are as follows:
  • We design fine-grained and computation-efficiency learnable masks that can capture malicious models’ characteristics in a uniform high-dimension space.
  • We develop SkyMask, a new attack-agnostic Byzantine-robust FL system, which is the first work applying fine-grained learnable masks to detect malicious clients and defend against Byzantine attacks.
  • We empirically evaluate SkyMask on various benchmarks under six SOTA attacking methods and compare them with existing defense methods, where experimental results show that SkyMask achieves up to 10% higher testing accuracy compared with existing defenses.
II Preliminaries & Related Work
II-A Federated Learning
A typical FL system includes $n$ clients and a server. Each client $i$ has a local dataset $D_{i}$, $i=1,\dots,n$. In each communication round $t$, the server selects a subset of the clients $\mathcal{N}_{t}$ to execute the following three steps: 1) Model distribution: The server distributes the global model $W_{t}$ to the selected clients. 2) Local training: The clients receive global model $W_{t}$ as the initial model $W^{i}_{t+1}$ for local training, and repeat $W_{t+1}^{i}:=W^{i}_{t+1}-\beta\nabla f(W^{i}_{t+1};D_{i})$ for $l$ local iterations, where $f(\cdot;\cdot)$ is the empirical loss function and $\beta$ denotes the local learning rate. The corresponding local model update is $\Delta W^{i}_{t+1}=W_{t}-W^{i}_{t+1}$, so for the server, it is equivalent to the local model $W^{i}_{t+1}$. 3) Model aggregation: After the training, each client $i$ uploads its model weights $W_{t+1}^{i}$ to the server for aggregation. FedAvg [1] performs weighted model averaging to update the global model: $W_{t+1}=\sum_{i\in\mathcal{N}_{t}}\frac{|D_{i}|}{\sum_{i\in\mathcal{N}_{t}}|D_{i}|}W_{t+1}^{i}$, where we set $|\mathcal{N}_{t}|=n$ for simplicity.
II-B Threat Model & Byzantine Attacks on FL
Malicious clients compromised by an attacker participate in the FL process and have access to the knowledge of neural network models, learning rates, and objective functions. Besides, the attacker can fully control these clients’ activities (e.g., modifying the model updates) and has complete access to their local datasets [4, 9, 12].
Existing Byzantine attacks on FL can be categorized into untargeted, targeted, or backdoor based on the adversary’s goal: 1) Untargeted attacks seek to corrupt the global model and minimize its accuracy on any test input [2, 4, 12, 5]. For example, Fang [4] adaptively trades off the attacking stealthiness and effectiveness at a fine granularity as an optimization problem. 2) Targeted attacks aim to minimize the accuracy on the specific test inputs, and maintain high accuracy on the rest [3, 14]. 3) Backdoor attacks reduce the utility of the global model on test inputs that contain a specific signal called trigger [15, 6].
II-C Byzantine-robust FL Algorithms
We classify existing representative Byzantine robust algorithms into two categories based on their strategies:
II-C1 Model-level defense strategies
FLDetector [16] determines whether a client is malicious by performing the consistency analysis on its model updates. Specifically, FLDetector predicts a client’s model update for each iteration based on historical records and compares it with the received model update. If they are inconsistent in multiple iterations, the client is probably malicious. FLTrust [9] trains a root model with a small clean root dataset at the server in each round, and measures the cosine similarity between a local model and the root model. The similarity metric dictates the weight in global model aggregation. Tolpegin [2] detects malicious participants by analyzing the characteristics of model updates. It standardizes all model updates, uses the principal component analysis (PCA) for dimension reduction, and identifies malicious clients via clustering algorithms. Krum [8] assumes the number of malicious clients is known as $n_{m}$. The server computes a score $s^{i}_{t+1}$ for each model update from the sum of squares of the distance between this model update and the $n-n_{m}-2$ nearest ones. Then, the one with the smallest score is selected as the global model update for the next round. Multi-Krum (MKrum) [8], a Krum variant, selects $c$ model updates with the lowest scores, and then aggregates them for a new global model.
II-C2 Greedy parameter-level filtering strategies
Trimmed-Mean (Trim) and Median [10] are coordinate-wise aggregation rules assuming the number of malicious clients $n_{m}$ is known, remove the parameters of the smallest and largest $n_{m}$ values, and average the remaining. Median retains the corresponding coordinate’s median value as the new global model’s value. Bulyan [11] also relies on a known $n_{m}$ and requires $n\geq 4n_{m}+3$ to ensure its robustness. It first iteratively applies MKrum to select $\theta$ local models, where $\theta<n-2n_{m}$, and uses a variant of Trim to aggregate the $\theta$ selected local models.
However, existing robust FL strategies can hardly defend against those fine-grained attacks without sacrificing training efficiency and model quality by utilizing coarse-grained similarity checking or greedy parameter-wise filtering, which motivates us to explore malicious client detection with fine-grained learnable masks.
II-D Motivating Learnable Masks
(a) The PCA of model updates
(b) The PCA of masks
Figure 1: Visualizing model updates and masks with PCA.
Binary masks have been widely used in deep learning and demonstrated the capability to capture neural network models’ intrinsic characteristics at the parameter-level. For example, Zhou et al[17] prune deep neural network models with binarized masks to improve training efficiency. FedMask [18] utilizes learnable binary masks to enable each device to learn a personalized and structured sparse neural network model that runs efficiently on devices.
We use a realistic experiment to show how learnable masks can be applied to detect malicious clients. In a swarm of 100 FL clients, we compromise 20 as malicious clients to apply Min-Sum attack. We train a four-layer CNN model on Fashion-MNIST dataset with the non-IID setting described in §IV-A1. We visualize the CNN model updates of all clients in one iteration using PCA in Figure 1(a), which shows how the fine-grained Min-Sum attack hides the malicious model updates among the benign updates. Hence, analyzing local model updates with PCA is insufficient to distinguish benign and malicious clients.
To detect such stealthy attacks, we apply fine-grained learnable masks that capture the parameter-level characteristics of local models. We aggregate all masked local models and train the masks together on a small root dataset, where the learnable masks learn to enhance the benign models and correct the malicious models (details in §III-B). Figure 1(b) shows the visualization of trained masks using PCA, where the masks applied to the malicious model updates are easily distinguished among all masks. This phenomenon reveals the capability of learnable masks that captures the difference between benign and malicious model updates.
II-E Root Dataset
Considering that the server has no root of trust to decide whether a model update is malicious or not, FLTrust [9] was the first to introduce the concept of root dataset. It can take advantage of the root dataset even if the distribution diverges from the overall data distribution or the size is less than one hundred. Therefore, it is easy for the server and FL manager to manually collect or generate a small root dataset. The rationality for using root datasets has been verified in several other approaches [19, 20].
III SkyMask’s Design
III-A Goals
We aim to design a server-based robust FL system. The server has a small representative root dataset following a previous work [9]. Moreover, this root dataset can be imbalanced without a major bias (analyzed in §IV-D1). We allow reasonable additional computation costs and extra storage space to defend against Byzantine attacks. We aim to detect all the malicious model updates and filter them out at each communication round to eliminate the impact of attacks. We should also keep as many benign clients as possible in the filtered FL client set to train a final global model with sufficient local data for satisfactory accuracy. For the proposed Byzantine-robust FL system, it should promise the following three features:
  • Robustness. The system should preserve the global model accuracy when attacked by malicious clients. In particular, we should maintain a high malicious client detection accuracy and a low misidentification rate of benign ones (i.e., false positive rate). The system should also stay effective on different datasets and model structures under different attacks.
  • Efficiency. The system should not introduce extra computation and communications overhead to the clients, which are often resource-constrained devices in FL. Besides, when performing malicious client detection and defense on the server, the system should only incur a proper computation complexity.
  • Scalability. The system should still support a large scale of clients and work under a high fraction of malicious clients (e.g., 80%) since FL systems usually involve a large number of client devices.
III-B SkyMask: Overview
When initializing the FL process, the server creates an initial global model and builds a root dataset as [9]. Figure 2 presents SkyMask’s workflow, where each communication round has six steps: Step 1, the server distributes the global model parameters to each client. Step 2, the clients load the parameters, train the local models and send back these model updates. Step 3, the server freezes received local models and applies learnable masks to each local model update. Step 4, all the masked local models are trained together on the root dataset to converge. Step 5, the server detects and removes the malicious clients by clustering all trained masks. Step 6, the server aggregates the remaining model updates into a new global model.
The key component of SkyMask is the detection mechanism based on learnable masks. Specifically, SkyMask constructs a same-size fine-grained mask for each local model update and trains these masks by optimizing the aggregation result of masked local models until they converge. Byzantine attacks take impairing global model performance as an optimization objective, so the poisoned model updates tend to make the global model far from normal behavior. Therefore, to correct potential misbehaviors of malicious models, the masks applied to the malicious models learn to mitigate the side effects of poisoned parameters through training on the root dataset. In this way, the learnable masks form a high-dimensional representation space to capture the characteristics of both malicious and benign model updates. Then, the clusters formed by masks are the basis for determining whether a client is malicious. After removing all the detected malicious model updates, SkyMask calculates the average of remaining model updates and uses it to update the global model.
III-C The Complete SkyMask Algorithm
Most existing defense strategies are based on the local model updates’ coarse-grained statistics or greedily filtering out outlier parameters. However, fine-grained attacks corrupt a small set of particular layers or parameters to work around existing defense strategies easily. Thus, we apply fine-grained learnable masks and clustering analysis to detect and defend against malicious clients. Our defense strategy has two stages: 1) mask initialization and training stage; 2) mask clustering and classification stage.
Figure 2: SkyMask’s workflow.
III-C1 Mask initialization and training
In the first stage, the server assigns a learnable mask $m_{i}$ of the same size as the model for each client $i$ and initializes them all with $1$s. The model aggregated from masked local models $\tilde{W}_{t+1}$ (called aggregated masked model) is computed by averaging the masked local models $m_{i}\odot W^{i}_{t+1}$. Then, the masks start training on the root dataset with frozen local models. At the beginning of iteration $t+1$, $\tilde{W}_{t+1}$ is calculated as:
$W^{i}_{t+1}=W_{t}-\Delta W^{i}_{t+1},$
$\tilde{W}_{t+1}=\sum^{n}_{i=1}\frac{1}{n}m_{i}\odot W^{i}_{t+1}.$ (1)
We train the aggregated masked model $\tilde{W}_{t+1}$ on the root dataset. Then, the update process back-propagates gradients with the help of $\tilde{W}_{t+1}$ to masks and applies a gradient descent algorithm to the masks.
Intuitively, we know that if we fix all masks as zeros except for one mask to be trained, there is also a solution for this optimization problem. If $\tilde{W}_{t+1}$ converges to the fixed optimal model $\hat{\tilde{W}}_{t+1}$, the problem is equivalent to solving a multivariate primary equation, which has infinite solutions. Moreover, suppose the mask value is not limited. In that case, some masks have very large values, and others are so small for the same dimension that these masks can not effectively characterize whether the local model is benign. So we cannot detect malicious clients with the help of this type of mask, and we should redesign the masks to show the contribution of each local model update in the optimized model. Binary can directly determine whether a parameter of model update participates in the global model, and binary masks have shown effectiveness in model pruning. So we use binary masks to extract the parameter-level characteristics of local models. If the malicious model parameters attempt to corrupt the aggregated model, the optimization process prefers to reject them and sets the corresponding parameters of the mask as 00.
III-C2 Approximation of binary masks
Binary parameters cannot be easily optimized by a gradient descent algorithm. Thus, we approximate the binary masks in training by setting the mask with real values and applying a sigmoid function $\sigma(\cdot)$ instead of a hard threshold to reduce the enormous gradient variance. Furthermore, it is easy for the gradient to be back-propagated to the real-value masks. The approximation of a binary mask is calculated as $\tilde{m}_{i}=\sigma(m_{i})$. Since the mask’s limitation is changed, the original Equation 1 is unsuitable. To adapt for the approximated binary masks, $\tilde{W}_{t+1}$ is re-written as:
$\tilde{W}_{t+1}=\frac{\sum^{n}_{i=1}\tilde{m}_{i}\odot W^{i}_{t+1}}{\sum^{n}_{i=1}\tilde{m}_{i}}.$ (2)
It can be formulated as an element-wise weighted averaging algorithm. If the parameters of a poisoned local model are toxic to the aggregated masked model $\tilde{W}_{t+1}$, the mask tends to reduce the involvement of these parameters to $\tilde{W}_{t+1}$ in the optimization phase. Thus, the 0-1 pattern in the mask can represent whether the corresponding local model is malicious.
The mask training task is not different from the main task in FL except for the variables. By using the stochastic gradient descent algorithm (SGD) to optimize these masks, the objective function is:
$f(\tilde{W}_{t+1},D_{r})=\sum\limits_{(x,y)\in D_{r}}L(\text{output}(x,\tilde{W}_{t+1}),y).$ (3)
During training the masks, masks are updated in each mask training iteration as follows:
$m_{i}:=m_{i}-\gamma\cdot\nabla_{m_{i}}f(\tilde{W}_{t+1},D_{r}),$ (4)
where $\gamma$ represents the mask learning rate.
This process will not stop until these masks converge. After the real-value masks converge, the final results of binary mask $\hat{m}_{i}$ are computed as Equation 5.
$\hat{m_{i}}[k]=\begin{cases}1,&\tilde{m}_{i}[k]>\tau\\ 0,&\tilde{m}_{i}[k]\leq\tau\end{cases},$ (5)
where $\hat{m}_{i}[k]$ represents the parameter located in $k$th dimension of binary mask $\hat{m}_{i}$ of client $i$ and $\tau$ is a threshold.
III-C3 Mask clustering and classification
These binary masks can sense the fine-grained poisoning attack and represent the parameter-level characteristics of each client’s model update. We use the PCA algorithm to reduce the dimensionality to $d$. Since the maximum dimensionality that the masks can be reduced to is limited by the number of masks $n$ and the dimensionality of masks $V$, we set $d=2$, which is the minimum possible number of clients in the FL system. Then, we apply the Gaussian mixture model for clustering and classification. If there are no attacks, the clustering result is only one cluster, and the server aggregates all the benign local models.
If the fraction of malicious clients is lower than 50%, the system can treat the larger cluster in two clusters as the benign client set. To cope with more general situations, SkyMask introduces a trusted root model $W^{r}_{t+1}$, which is initialized by $W_{t}$ and “locally” trained $l$ iterations on the root dataset as other clients do:
$W^{r}_{t+1}:=W^{r}_{t+1}-\beta\nabla f(W_{t},D_{r}).$ (6)
Then the server lets it join in the model set before the mask training stage, assigns one mask for it and trains this mask together with other masks. When all the masks converge, the clients represented by the masks in the same cluster as the mask corresponding to the trusted root model are considered benign. This way, SkyMask can handle various malicious fractions from zero to very high. In order not to interfere with the global model convergence, we should emphasize that the trusted root model is only used for malicious model detection and will not be included in the benign model aggregation phase.
SkyMask plays $R$ rounds for server-clients communication, and each round contains six steps, as §III-B describes. The critical point of SkyMask is our fine-grained malicious client detection strategy mentioned above, which can provide the remaining benign client set $U_{b}$ for aggregating a new global model. The global model for the next communication round $W_{t+1}$ is calculated as:
$W_{t+1}=W_{t}-\alpha\cdot\sum\limits_{i\in U_{b}}\frac{|D_{i}|}{\sum_{i\in U_{b}}|D_{i}|}\Delta W^{i}_{t+1},$ (7)
where $\alpha$ denotes the global learning rate. It should be noted that SkyMask is more concerned with malicious client detection and can be added as an additional module to the aggregation algorithm. Therefore, it can even enhance the existing Byzantine-robust aggregation algorithms. If it completes the malicious detection task, removing only the malicious model updates and leaving the benign model updates untamed, its convergence will be consistent with the original aggregation algorithm.
Figure 3: Visualizing the training process of masks applied to malicious and benign models, respectively.
III-D Why SkyMask Works
We conduct an example experiment to show how SkyMask works. We choose a four-layer CNN model and Fashion-MNIST dataset and set 20% of 100 clients as malicious. Under Fang-Trim attack, we take the mask training process in one server-client communication round. We calculate the proportion of “1”s in each part of the binary mask, and a smaller value means fewer parameters of this client are selected for the aggregated masked model. For simplicity, here only shows 20 masks for malicious and 20 for benign clients in each mask training iteration.
As Figure 3 shown, during the mask training process, the binary parameters in the “conv1.weight”, “conv2.weight” and “fc1.weight” parts of malicious model updates are set to zero more than those of benign updates. It is related to the fact that the attack prefers to attack the parameters in specific parts. The masks gradually reduce the poisoning effect of these toxic parameters during the training so that we can classify malicious clients with the help of these masks. Therefore, SkyMask can perceive the poisoning attacks at the parameter-level in the model updates without knowing the details of the attack, which makes it attack-agnostic.
III-E Complexity
Let $O(T)$ denote the computational complexity of one local training iteration, and the model parameters’ dimensionality is $V$. To update the masks once, the server calculates the gradient of each mask from the gradient of the aggregated masked model $\tilde{W}_{t+1}$ by backpropagation, and its computational complexity is $O(nV)$.
Hence, the total computational complexity is $O((T+nV)t_{m})+O(nV)=O(Tt_{m}+nVt_{m})$, where $t_{m}$ is the number of mask training convergence iterations. Since there is no data dependency between any mask parameters, we can use parallel training [21] to accelerate the mask optimization process so that the time overhead can be further reduced, trading computation power for time.
FLTrust only needs to calculate the similarity of local models and its baseline model, whose computational complexity is $O(nV)$. Trim sorts each parameter of local models and choose the middle part of the values, so its computational complexity is $O(nV\log n)$. Krum calculates the distance score of each local model, which is the sum of the distance between one local model and its $n-n_{m}-2$ nearest local models, so its computational complexity is $O(n^{2}V)$.
The above analysis shows that the additional cost of the mask learning mechanism over other defense methods is equivalent to some iterations of local training, which are acceptable compared to the computational power of the server.
The space complexity of existing FL defense algorithms is either $O(n+nV)$ or $O(n^{2}+nV)$. Since we assign a mask for each client, the additional space complexity is also proportional to the number of clients $n$. It is proportional to the space complexity of the global model parameters because the masks have the same dimensionality as the global model parameters. Considering that the server stores $n$ local models’ parameters in the current round, the total space cost is about twice the others, and its space complexity is $O(nV+nV)=O(nV)$, which is reasonable.
IV Evaluation
We implement a prototype of SkyMask based on PyTorch and evaluate our algorithm with existing Byzantine-robust FL aggregation algorithms and two malicious client detection algorithms. To verify the generalization of our method, we conduct experiments using several popular datasets and pair them with different models, respectively. We evaluate our approach with six popular attacks, explore the impact of the fraction of malicious clients, and verify its scalability. In sensitivity analysis, we discuss the influence of binarization threshold value, the number of features after dimensionality reduction, and the root dataset’s data distribution on our detection method’s performance. We repeat each experiment three times and average the results. All experiments run on a cloud server with eight NVIDIA RTX 2080 GPUs.
On Fashion-MNIST and HAR datasets, SkyMask improves the testing accuracy by 1–3% compared with the best result of all five Byzantine-robust algorithms under the same attacks. More significantly, our method gains a 10% improvement in testing accuracy on CIFAR-10 dataset under Min-Max attack.
IV-A Experiment Settings
IV-A1 IID and non-IID Datasets
Following previous work [4, 9], in different datasets for classification tasks, if there are $C$ classes, we construct $C$ groups. The probability of a training sample with label $k$ assigned to group $k$ is $q>0$ and to other group is $\frac{1-q}{C-1}$. Each client randomly chooses one group and samples a certain number of samples uniformly from this group for its local dataset. The probability $q$ controls the degree of non-IID. If $q=\frac{1}{C}$, the dataset for each client is IID. For the root dataset, we randomly sample a certain number of samples from the training dataset with a bias probability $p$ whose meaning is similar to $q$ and then split the rest for each client in the way mentioned above. We use Fashion-MNIST, CIFAR-10, and HAR dataset [22] as the evaluation datasets, which have been widely used in prior FL studies. The Fashion-MNIST dataset consists of a 60K image samples training set and a 10K samples test set, where each image is associated with a label from 10 classes. The CIFAR-10 dataset composes 50K training images and 10K test images in 10 classes. They are colorful with three channels, contain many features noise, and have varying object scales. For both Fashion-MNIST and CIFAR-10, we use $q=0.5$ to conduct the non-IID experiments as [4]. By setting $q=0.1$, we also conduct the IID experiments as [12].
HAR dataset is built by recording 30 participants’ daily living with sensors, consisting of 7,352 pieces of 561-feature vectors with time and frequency domain variables. The main task is to classify the activity into one of six human activities. Since they are collected naturally by 30 devices, it is unnecessary to divide the dataset manually. We randomly assign 70% of each device’s dataset as the training dataset and 30% to be the testing data.
IV-A2 Models
We construct a CNN with two convolutional layers and two linear layers as the global model for Fashion-MNIST dataset. For CIFAR-10 dataset, to verify SkyMask’s effectiveness on large models in FL, we use a more complex model, ResNet20. We select a linear regression classifier for HAR dataset.
IV-A3 Baseline attacks
The experiments cover untargeted, targeted, and backdoor attacks, including the following baselines:
  • Fang attack [4]: Fang attack maximizes the deviation of the global model towards the opposite direction of the unattacked global model update. By using Trim and Krum as the aggregation algorithm in the attack, we obtain two different attacks: Fang-Trim and Fang-Krum.
  • Label-flipped (LF) [2]: Following the same setting as Fang [4], we flip label $k$ to $C-k-1$ for each sample on malicious clients, where $k\in\{0,\dots,C-1\}$ and $C$ denotes the total number of dataset classes.
  • AGR-agnostic [12], i.e., Aggregation algorithm agnostic attack, aims to limit the range of malicious updates and maximize the perturbation. It has two variants: Min-Max attack restricts the range based on the maximum value of the L-$2$ distance between any pair of benign model updates. Min-Sum attack is based on the minimum value of the sum of squares of the distances between any benign model update and all other benign model updates.
  • Scaling attack [15]: Scaling attack is a targeted backdoor attack. It poisons the local dataset with trigger-embedded samples and scales up the model updates before sending them to the server. We use the same pattern trigger in LIE [5] and follow the same data augmentation scheme as FLTrust [9].
IV-A4 Baseline defenses
In §IV-B, to show the defense effectiveness, we compare SkyMask with five Byzantine-robust FL aggregation algorithms: FLTrust [9], Trim [10], Krum, MKrum [8] and Bulyan [11]. In §IV-C, to reveal the significance of learnable masks in malicious client detection, we compare SkyMask with two malicious client detection algorithms: Tolpegin defense [2] and FLDetector [16].
TABLE I: FL testing accuracy under different attacks and attack success rates of Scaling attack. The experimental results of Scaling attack is in the form of “testing accuracy / attack success rate”.
Dataset (Model) Attack FedAvg FLTrust Trim Krum MKrum Bulyan SkyMask-NR SkyMask
Fashion -MNIST (CNN) non-IID None 0.89 0.89 0.88 0.83 0.88 0.87 0.89 0.89
LF 0.84 0.86 0.84 0.83 0.89 0.87 0.89 0.89
Min-Max 0.58 0.89 0.70 0.83 0.64 0.65 0.89 0.89
Min-Sum 0.80 0.89 0.73 0.47 0.82 0.75 0.89 0.89
Fang-Trim 0.42 0.89 0.67 0.82 0.75 0.85 0.89 0.89
Fang-Krum 0.86 0.89 0.84 0.47 0.70 0.78 0.89 0.89
Scaling 0.80 / 0.21 0.89 / 0.08 0.87 / 0.13 0.82 / 0.09 0.88 / 0.10 0.87 / 0.10 0.89 / 0.10 0.89 / 0.10
Fashion -MNIST (CNN) IID None 0.90 0.89 0.89 0.86 0.89 0.87 0.89 0.90
LF 0.81 0.88 0.84 0.86 0.89 0.87 0.89 0.90
Min-Max 0.63 0.89 0.76 0.86 0.59 0.72 0.90 0.90
Min-Sum 0.85 0.89 0.78 0.55 0.84 0.68 0.90 0.90
Fang-Trim 0.53 0.89 0.74 0.86 0.84 0.86 0.90 0.90
Fang-Krum 0.88 0.89 0.86 0.26 0.71 0.78 0.90 0.90
Scaling 0.83 / 0.21 0.89 / 0.10 0.88 / 0.14 0.85 / 0.11 0.89 / 0.11 0.87 / 0.11 0.89 / 0.10 0.90 / 0.10
CIFAR-10 (ResNet20) non-IID None 0.77 0.75 0.77 0.54 0.72 0.72 0.76 0.76
LF 0.71 0.74 0.71 0.55 0.69 0.68 0.75 0.76
Min-Max 0.58 0.68 0.70 0.52 0.63 0.55 0.77 0.77
Min-Sum 0.66 0.72 0.74 0.32 0.66 0.68 0.75 0.77
Fang-Trim 0.10 0.68 0.19 0.52 0.53 0.49 0.75 0.76
Fang-Krum 0.58 0.75 0.48 0.19 0.36 0.40 0.77 0.77
Scaling 0.10 / 1.00 0.74 / 0.13 0.72 / 0.51 0.53 / 0.08 0.73 / 0.10 0.73 / 0.10 0.77 / 0.09 0.77 / 0.10
CIFAR-10 (ResNet20) IID None 0.79 0.78 0.78 0.63 0.73 0.73 0.79 0.79
LF 0.76 0.76 0.74 0.64 0.72 0.73 0.77 0.78
Min-Max 0.68 0.61 0.58 0.56 0.65 0.69 0.79 0.79
Min-Sum 0.73 0.75 0.62 0.30 0.66 0.70 0.79 0.79
Fang-Trim 0.10 0.70 0.38 0.62 0.54 0.69 0.78 0.78
Fang-Krum 0.68 0.77 0.60 0.42 0.34 0.34 0.79 0.79
Scaling 0.10 / 1.00 0.76 / 0.12 0.75 / 0.67 0.62 / 0.10 0.74 / 0.10 0.74 / 0.10 0.78 / 0.11 0.79 / 0.10
HAR (LR) None 0.96 0.96 0.96 0.81 0.95 0.94 0.96 0.96
LF 0.88 0.95 0.81 0.86 0.79 0.74 0.96 0.96
Min-Max 0.89 0.95 0.93 0.76 0.83 0.66 0.96 0.96
Min-Sum 0.96 0.95 0.94 0.30 0.86 0.83 0.96 0.96
Fang-Trim 0.23 0.96 0.80 0.77 0.65 0.82 0.96 0.96
Fang-Krum 0.96 0.96 0.93 0.15 0.75 0.71 0.96 0.96
Scaling 0.72 / 0.20 0.96 / 0.20 0.92 / 0.16 0.84 / 0.19 0.96 / 0.20 0.94 / 0.19 0.96 / 0.19 0.96 / 0.20
IV-A5 FL parameter settings
We use 100 clients in Fashion-MNIST and CIFAR-10 experiments, and in HAR, we use 30 clients, following previous study [9]. To boost the effectiveness of attacks, we apply multiple-client attacks instead of single-client attacks, and the default fraction of malicious clients is set to 20%. The default binarization threshold $\tau$ is $0.5$, and the root dataset bias probability $p=0.1$.
We set both global model learning rate $\alpha$ and local model learning rate $\beta$ to $0.5$ in all experiments. For mask learning rate $\gamma$, we set $1\times 10^{7}$ in Fashion-MNIST and HAR datasets but $1\times 10^{8}$ in CIFAR-10 dataset. This is because the order of magnitude of the gradient is $10^{-8}-10^{-7}$, which declines along the model’s dimensionality increases, so we should set it larger to help masks converge. In CIFAR-10, we allow the system to execute server-client communication for $500$ rounds and let clients do $l=5$ local training iterations per round. While in Fashion-MNIST and HAR, the system runs $2,500$ and $1,000$ communication rounds, respectively, and each client only does a single local training iteration per round.
In our experiments, the root dataset has 100 samples and no intersection with the training and test datasets. This setting is similar to FLTrust [9]. 100 is enough for our defense, and it is easy for the server to collect. So we do not discuss the influence of the size of the root dataset.
IV-A6 SkyMask variants
Intuitively, if the fraction of malicious clients is less than 50% in the real world, we can choose the larger one in two clusters of masks as benign. This way, we can further reduce the computational pressure on the server. To verify the effectiveness with or without the trusted root model, we consider the following two schemes: SkyMask-NR (no trusted root model): The FL system does not generate a trusted root model, and after all the masks are converged, the system chooses the larger one in two clusters of the masks as benign for aggregation. SkyMask: The FL system uses the defense strategy fully following the default steps described in §III-C.
IV-A7 Metrics
Our experiments compare the testing accuracy of global models of various FL defenses under different attacks, obtained from the prediction accuracy of the global model on the test dataset. The attack success rate under Scaling attack is the fraction of the samples in the test dataset that are identified as the target label after being embedded with triggers.
To show the malicious client detection ability, we also record false positive rate (FPR) and false negative rate (FNR) in the malicious client detection stage. FPR indicates the number of benign clients that are misidentified as malicious clients divided by the total number of benign clients, and FNR indicates the number of malicious clients that are misidentified as benign clients divided by the total number of malicious clients. The Mean FPR/FNR is the average of FPR/FNRs in each communication round.
IV-B The Defense Effectiveness of SkyMask
We first certificate the normal defensive capability under a low fraction of attacks. Then, there is also possibly no attacker in the FL system, so the defense methods should not affect the global model’s performance. Eventually, we consider an extreme condition that the fraction of malicious clients is high.
Figure 4: The impact of high fractions of malicious clients.
IV-B1 Under a low fraction of attacks
By default, we set the fraction of malicious clients to 20% following the default setting in [9]. Table I shows that, on Fashion-MNIST dataset, Multi-Krum and Bulyan algorithms perform robustness under LF attack, but they fail in the other attacks. FLTrust is relatively good that in a few experiments it achieves the same performance as ours in the non-IID case, while under LF attack, it has a 3% accuracy loss. Our SkyMask achieves the highest accuracy rate on Fashion-MNIST under different attacks compared to all others. Scaling attack does not show an effective impact on all the defense methods.
To demonstrate the generalization, we conduct the experiments on CIFAR-10 dataset using a larger model, ResNet20. In the IID and non-IID cases, FLTrust does not perform as well as it does on FashionMNIST dataset, with a severe drop in accuracy under multiple attacks (especially the Min-Max attack). The other defense methods have little or no defensive effect. On the contrary, our SkyMask reaches the highest accuracy under attacks. In the IID case, our method achieves the most significant gap with other algorithms in Min-Max attack experiments. The testing accuracy is up to $0.79$, while the accuracy results obtained by others are less than $0.69$.
Since the samples with target label 0 are 10% of the test dataset in FashionMNIST and CIFAR-10, the attack success rate is about 10% if there is no attack. We can see that FedAvg is overwhelmed and Trim suffers from the Scaling attack severely on CIFAR-10. SkyMask still keeps the robustness and performs well.
The HAR dataset has a real-world data distribution, making the experiment more meaningful and closer to reality. We can see in the HAR part of Table I that Trim, Krum, and their variants cannot work well. FLTrust always has a little gap with the accuracy of models not under attack, which means it could not eliminate the impact of attacks, but it is still acceptable. Unlike all these defense methods, our SkyMask keeps the global model immune to attacks.
Due to the attack designers’ optimization for the abnormality degree, only some particular parameters of the poisoning model updates changed with emphasis. It can be seen that the four fine-grained model poisoning attacks (Min-max, Min-Sum, Fang-Trim, and Fang-Krum) bypass the existing defense methods in many cases.
In all these experiments, SkyMask not only achieves the highest testing accuracy but reaches the same accuracy level as unattacked FedAvg’s, i.e., the same accuracy or no more than 1% accuracy loss.
As Section IV-A6 describes, the server does not need to consume resources to generate a root model if the fraction of malicious clients is less than 50%. In the 20% case, as shown in Table I, SkyMask-NR achieves good results under any attack and on any dataset, which shows the same defensive capability as SkyMask. Therefore, if the server confirms that the number of malicious clients is less than half the total number, it can choose SkyMask-NR as the defense strategy to reduce the overhead while maintaining the same defensive capability. For generalization, SkyMask can be applied to handle more cases.
IV-B2 Under no attack
Some attacks send poisoned models to the server only at certain communication rounds, and some model updates are often mistaken for poisoning by existing defense methods due to data heterogeneity. The system collapses if the global model performance suffers from the defense strategy, even when there is no attack. So when there is no attack, the expected result is that any method should not affect the performance of the basic aggregation algorithm FedAvg can achieve.
In Table I, almost all other robust aggregation algorithms affect the convergence prediction accuracy in the no-attack case. For instance, FLTrust has an accuracy loss of about 2% for the non-IID case on CIFAR-10 dataset, and Krum algorithm has an accuracy loss ranging from about 4% to 20% on different datasets.
At the same time, our SkyMask maintains a comparable performance as unattacked FedAvg on all datasets with a difference of less than 1%. The existing robust aggregation algorithms either choose a subset of model updates that seem benign or use model updates’ statistics to correct the impact. Hence, it is possible to introduce harmful impacts without attack. If there is no attack, our masks form only one cluster, or there are only several outliers. So the server selects most clients and has the same or very similar convergence with FedAvg in both IID and non-IID cases.
TABLE II: Testing accuracy, FPR and FNR of different malicious client detection methods under different attacks. The experimental results of Scaling attack is in the form of “testing accuracy / attack success rate”.
Dataset (Model) Attack Testing accuracy FPR FNR
Tolpegin FLDetector SkyMask Tolpegin FLDetector SkyMask Tolpegin FLDetector SkyMask
Fashion -MNIST (CNN) non-IID None 0.89 0.86 0.89 / / / / / /
LF 0.89 0.86 0.89 0.11% 0.05% 0.00% 0.38% 0.00% 0.02%
Min-Max 0.88 0.08 0.89 0.38% 96.94% 0.27% 0.80% 100% 0.40%
Min-Sum 0.64 0.38 0.89 41.1% 100% 0.26% 84.8% 100% 0.00%
Fang-Trim 0.88 0.86 0.89 0.00% 0.00% 0.00% 0.00% 0.00% 0.00%
Fang-Krum 0.85 0.10 0.89 28.2% 99.46% 0.20% 60.4% 100% 0.00%
Scaling 0.89 / 0.10 0.89 / 0.10 0.89 / 0.10 0.00% 0.00% 0.00% 0.03% 0.00% 0.00%
Fashion -MNIST (CNN) IID None 0.90 0.88 0.90 / / / / / /
LF 0.89 0.87 0.90 0.07% 0.02% 0.00% 0.00% 0.00% 0.00%
Min-Max 0.89 0.87 0.90 0.06% 0.00% 0.00% 0.00% 0.81% 0.00%
Min-Sum 0.88 0.49 0.90 15.34% 100% 0.01% 7.60% 100% 0.00%
Fang-Trim 0.90 0.88 0.90 0.00% 0.00% 0.00% 0.00% 0.00% 0.00%
Fang-Krum 0.89 0.13 0.90 8.04% 99.60% 0.19% 6.40% 100% 0.00%
Scaling 0.89 / 0.11 0.90 / 0.11 0.90 / 0.10 0.00% 0.00% 0.00% 0.68% 0.00% 0.00%
CIFAR10 (ResNet20) non-IID None 0.76 0.75 0.76 / / / / / /
LF 0.70 0.72 0.76 13.1% 0.03% 4.72% 19.9% 25.53% 2.60%
Min-Max 0.61 0.11 0.77 36.5% 100% 0.00% 88.0% 100% 0.00%
Min-Sum 0.59 0.31 0.77 38.8% 100% 0.00% 78.0% 100% 0.00%
Fang-Trim 0.76 0.74 0.76 0.00% 0.03% 0.00% 0.00% 0.00% 0.00%
Fang-Krum 0.17 0.31 0.77 37.4% 87.18% 0.00% 84.0% 100% 0.00%
Scaling 0.76 / 0.10 0.74 / 0.10 0.77 / 0.11 0.00% 0.00% 0.00% 0.10% 0.00% 0.00%
CIFAR10 (ResNet20) IID None 0.75 0.79 0.79 / / / / / /
LF 0.73 0.75 0.78 3.45% 0.00% 3.24% 5.80% 4.26% 0.00%
Min-Max 0.74 0.25 0.79 0.05% 100% 0.00% 0.00% 100% 0.00%
Min-Sum 0.74 0.31 0.79 0.14% 100% 0.00% 1.00% 100% 0.00%
Fang-Trim 0.74 0.77 0.78 0.03% 0.00% 0.00% 0.00% 0.05% 0.00%
Fang-Krum 0.74 0.18 0.79 4.04% 94.60% 0.00% 10.60% 100% 0.00%
Scaling 0.79 / 0.09 0.78 / 0.10 0.79 / 0.10 0.00% 0.00% 0.00% 0.00% 0.00% 0.00%
HAR (LR) None 0.96 0.92 0.96 / / / / / /
LF 0.96 0.68 0.96 0.22% 13.27% 0.00% 0.13% 100% 0.00%
Min-Max 0.96 0.07 0.96 1.38% 100% 0.83% 2.00% 100% 0.00%
Min-Sum 0.95 0.53 0.96 31.6% 100% 4.53% 32.0% 100% 4.00%
Fang-Trim 0.95 0.96 0.96 0.00% 0.00% 0.96% 0.00% 0.00% 0.00%
Fang-Krum 0.95 0.13 0.96 26.4% 100% 0.03% 29.3% 100% 0.00%
Scaling 0.96 / 0.20 0.95 / 0.19 0.96 / 0.20 0.25% 0.00% 0.00% 12.58% 0.00% 0.08%
IV-B3 Under a high fraction of attacks
The experiments mentioned in the previous sections set the fraction of malicious clients to 20%, while in a real application environment, an attacker can control a larger fraction of malicious clients for the attack. Therefore, to verify the defensive capability of our system in this case, we use CNN as the global model on Fashion-MNIST dataset in the non-IID case and conduct comparison experiments with 40%, 60%, and 80% of malicious clients. Since Bulyan requires that the number of malicious clients $m$ should satisfy $n\geq 4m+3$, these experiments do not include this algorithm. Trim requires that the number of malicious clients should meet $n\geq 2m+1$, and Multi-Krum requires that $n-c>2m+2$, so these aggregation algorithms are only included in 40% experiments.
The experimental results are shown in Figure 4. Most algorithms perform well under Scaling attack except for FedAvg and Krum. When the fraction of malicious clients is 40%, Trim loses its defensive capability completely, and Multi-Krum can only work under LF attack. Krum can reduce the impact of attacks under LF and Fang-Trim attacks, but the accuracy loss is still large. Although the accuracy of FLTrust fluctuates in some communication rounds, it still tends to converge to an unattacked level. From Figure 4(c), Figure 4(d) and Figure 4(e), we can see that SkyMask converges more stable than FLTrust when they are attacked by Fang-Krum attack, Min-Max attack and Min-Sum attack. In Figure 4(a) and Figure 4(e), the accuracy obtained by SkyMask is 2% to 4% higher than FLTrust’s under LF and Min-Sum attacks.
When the fraction is more than 50%, only FLTrust and SkyMask maintain the defensive capability, but FLTrust’s performance worsens as the fraction rises. FLTrust’s testing accuracy fluctuates even more, and the global model even converges in the wrong direction (e.g., in Figure 4(q), the accuracy obtained by FLTrust decreases from 85% to 81% after a sharp fluctuation). Under other attacks, the increasing fraction influences FLTrust’s convergence speed, and SkyMask converges faster than it.
(a) Fang-Krum attack, Fashion-MNIST
(b) Min-Sum attack, Fashion-MNIST
(c) LF attack, CIFAR-10
Figure 5: The impact of detection failure and the training efficiency of Tolpegin defense, FLDetector and SkyMask.
(a) On FashionMNIST Dataset (non-IID) with CNN
(b) On FashionMNIST Dataset (IID) with CNN
(c) On CIFAR-10 Dataset (non-IID) with ResNet20
(d) On CIFAR-10 Dataset (IID) with ResNet20
Figure 6: The impact of root data distribution on SkyMask.
IV-C The Significance of Learnable Masks
We conduct comparative experiments on SkyMask and the other malicious client detection algorithms, i.e., Tolpegin defense [2] and FLDetector [16] (combined with Median). We sample the result of malicious client detection every ten communication rounds and calculate all the metrics described in §IV-A7.
In Table II, we can see that FLDetector only works under LF attack, Fang-Trim attack and Scaling attack. Min-Max and Min-Sum attacks make FLDetector completely confuse the malicious model updates with the benign model updates.
In the Fashion-MNIST part of Table II, we can see that it is hard for Tolpegin defense to detect malicious clients stably, especially under Fang-Krum attack and Min-Sum attack. Its FPR and FNR are pretty high, so it fails to defend against attacks effectively and loses a lot of benign information. Quite the opposite, SkyMask perfectly completes the detection task and achieves an FNR of less than 4‰  on Fashion-MNIST dataset with a CNN model, which means that the server hardly misses any malicious client, and it can eliminate the harm of attacks. SkyMask also achieves an FPR of less than 3‰ , so it can preserve most benign clients and better characterize the global data distribution.
Figure 5(a) and Figure 5(b) show the trend of testing accuracy on non-IID Fashion-MNIST with a CNN global model under Fang-Krum attack and Min-Sum attack. Detection failure means that the FL defense system does not detect all the malicious clients in that round. We can see that SkyMask gradually rises and converges without detection failure. On the contrary, Tolpegin defense converges but with poor testing accuracy under Fang-Krum attack, and it is damaged under Min-Sum attack. Each detection failure causes a fluctuation in testing accuracy.
In the CIFAR-10 part of Table II, Tolpegin defense performs even worse in the non-IID case with large model architecture (ResNet20). Except for Fang-Trim attack, all other attacks greatly harm it. The large FNR of Tolpegin defense means many malicious clients complete their attacks. The large FPR of Tolpegin defense shows that the remaining benign clients may lose some representation, so Tolpegin defense achieves a lousy performance. In the IID case, Tolpegin defense resumes partial defensive capability, but there is still about 10% FNR under Fang-Krum attack. However, our SkyMask achieves the same prediction accuracy as a no-attack case, with FPR less than 5% and FNR less than 3%.
LF attack on CIFAR-10 dataset is more stealthy, and the FNR of SkyMask rises to 2.6%. As Figure 5(c) shows, at the beginning of FL training under LF attack, determining whether the difference between clients comes from data distribution or malicious client’s anomaly operation is more challenging. However, this does not matter since the LF attack is weak at the beginning of the training process. The global model still moves towards convergence; after several communication rounds, SkyMask can resume normal detection ability and identify these malicious clients. This mistake does not prevent our method from helping the global model reach convergence and high testing accuracy.
TABLE III: The impact of threshold in binarization on SkyMask.
Dataset (Model) Attack Acc. FPR FNR
0.3 0.4 0.5 0.6 0.7 0.3 0.4 0.5 0.6 0.7 0.3 0.4 0.5 0.6 0.7
Fashion -MNIST (CNN) non-IID None 0.89 0.89 0.89 0.89 0.89 / / / / / / / / / /
LF 0.88 0.88 0.89 0.89 0.88 0.01% 0.02% 0.00% 0.03% 0.32% 0.08% 0.02% 0.02% 0.00% 0.04%
Min-Max 0.89 0.89 0.89 0.89 0.89 0.00% 0.00% 0.27% 0.00% 0.00% 0.00% 0.00% 0.40% 0.00% 0.00%
Min-Sum 0.88 0.89 0.89 0.89 0.88 0.04% 0.20% 0.26% 0.34% 0.89% 0.00% 0.40% 0.00% 0.00% 1.60%
Fang-Trim 0.88 0.89 0.89 0.88 0.88 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00%
Fang-Krum 0.89 0.89 0.89 0.89 0.89 0.18% 0.28% 0.20% 0.40% 0.28% 0.40% 0.00% 0.00% 0.40% 0.00%
Scaling 0.89 / 0.10 0.89 / 0.12 0.89 / 0.10 0.89 / 0.10 0.89 / 0.11 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00%
Fashion -MNIST (CNN) IID None 0.88 0.89 0.90 0.89 0.89 / / / / / / / / / /
LF 0.89 0.89 0.90 0.90 0.89 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00%
Min-Max 0.89 0.90 0.90 0.90 0.89 0.08% 0.00% 0.00% 0.00% 0.01% 0.00% 0.00% 0.00% 0.00% 0.00%
Min-Sum 0.89 0.89 0.90 0.90 0.89 0.04% 0.05% 0.01% 0.01% 0.34% 0.13% 0.00% 0.00% 0.00% 1.00%
Fang-Trim 0.89 0.90 0.90 0.90 0.90 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00%
Fang-Krum 0.89 0.89 0.90 0.90 0.89 0.77% 0.56% 0.19% 0.05% 0.07% 1.00% 0.20% 0.00% 0.00% 0.20%
Scaling 0.89 / 0.12 0.90 / 0.11 0.90 / 0.10 0.90 / 0.11 0.90 / 0.10 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00%
CIFAR-10 (ResNet20) non-IID None 0.76 0.76 0.76 0.76 0.76 / / / / / / / / / /
LF 0.76 0.76 0.76 0.76 0.76 4.72% 4.33% 2.03% 2.40% 6.65% 5.70% 2.80% 2.60% 4.00% 6.30%
Min-Max 0.77 0.77 0.77 0.77 0.77 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00%
Min-Sum 0.77 0.77 0.77 0.77 0.77 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00%
Fang-Trim 0.76 0.77 0.76 0.76 0.76 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00%
Fang-Krum 0.77 0.77 0.77 0.77 0.77 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00%
Scaling 0.76 / 0.09 0.76 / 0.12 0.77 / 0.12 0.77 / 0.10 0.77 / 0.10 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00%
CIFAR-10 (ResNet20) IID None 0.79 0.78 0.79 0.78 0.78 / / / / / / / / / /
LF 0.76 0.78 0.78 0.79 0.78 5.31% 3.40% 3.24% 1.25% 8.04% 5.45% 1.75% 0.00% 4.45% 8.45%
Min-Max 0.78 0.79 0.79 0.79 0.78 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00%
Min-Sum 0.78 0.79 0.79 0.79 0.79 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00%
Fang-Trim 0.78 0.79 0.78 0.78 0.78 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00%
Fang-Krum 0.78 0.79 0.79 0.78 0.78 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00%
Scaling 0.79 / 0.11 0.79 / 0.10 0.79 / 0.10 0.79 / 0.09 0.79 / 0.11 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00%
HAR (LR) None 0.96 0.96 0.96 0.96 0.96 / / / / / / / / / /
LF 0.96 0.96 0.96 0.96 0.96 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00%
Min-Max 0.96 0.96 0.96 0.96 0.96 1.70% 1.57% 0.97% 2.50% 0.43% 1.00% 0.00% 1.00% 1.00% 1.40%
Min-Sum 0.96 0.96 0.96 0.96 0.96 2.17% 2.63% 0.87% 3.60% 5.70% 2.50% 2.00% 0.70% 1.00% 3.00%
Fang-Trim 0.96 0.96 0.96 0.96 0.96 0.16% 0.00% 0.00% 0.00% 0.00% 2.63% 0.00% 0.00% 0.00% 0.00%
Fang-Krum 0.96 0.96 0.96 0.96 0.96 0.70% 0.60% 0.00% 0.07% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00%
Scaling 0.96 / 0.22 0.96 / 0.19 0.96 / 0.20 0.96 / 0.16 0.96 / 0.20 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00%
Figure 7: The impact of root data distribution on model convergence.
For HAR dataset, Tolpegin defense’s testing accuracy seems good even when FNR rises to about 30%. The reason is that since the classification task is simple, even if detection failure in some rounds leads to the fluctuation of the testing accuracy, the testing accuracy can still rise to a reasonable level through successive rounds of detection success. In contrast, SkyMask achieves a good malicious client detection result with FNR less than 4%, maintaining high testing accuracy.
IV-D Sensitivity analysis
(a) On FashionMNIST Dataset (non-IID) with CNN
(b) On FashionMNIST Dataset (IID) with CNN
(c) On CIFAR-10 Dataset (non-IID) with ResNet20
(d) On CIFAR-10 Dataset (IID) with ResNet20
(e) On HAR Dataset with LR
Figure 8: The impact of dimensionality reduction on SkyMask.
IV-D1 The Impact of Root Dataset Distribution
The server’s root dataset is randomly sampled from the global data distribution by default. It is most likely to collect a root dataset with bias in the real world. However, we can prove that as soon as the bias is not too large ($p\leq 0.8$), our method can still take advantage of the root dataset and complete the malicious client detection task. We experiment on Fashion-MNIST with a CNN global model and CIFAR-10 with a Resnet20 model. The results are as testing accuracy part of Figure 6 shown: most of the testing results of the global model obtained by our method are very close or equal to the model obtained by FedAvg under no attack. However, in some cases, such as Figure 6(b) under Min-Max attack, and Figure 6(d) under no attack and LF attack, the testing accuracy is not acceptable when the bias probability of the root dataset is 0.95. Considering that the root dataset has a total of 100 samples, the lack of some classes of samples in the root dataset is possible when the bias probability of the root dataset is 0.95, so this result is reasonable.
The FPR and FNR results in the figures show that the malicious client detection accuracy of SkyMask fluctuates as the degree of non-IID of the root data changes but remains very low. Our method does not fail until the bias probability of the root dataset exceeds 80%. As long as the generated root dataset has a bias not so significant, SkyMask can complete its malicious client detection task and keep the global model a good performance.
In Figure 7, we show the variation curve of accuracy of SkyMask for non-IID FashionMNIST experiments. The convergence speed reduces along with the data distribution bias increasing, especially under Fang-Trim attack and Scaling attack. Although the root dataset’s different bias probability ($p\leq 0.8$) does not affect the convergence result, this phenomenon tells us that the imbalanced root dataset can slow down the global model’s convergence, and we had better choose a good root dataset for the server.
IV-D2 The Impact of Thresholds in Binarization
As the description of the processing of binary masks in Section III-C, the server should choose a threshold $\tau$. Our optimization strategy makes the parameters in masks converge to 0 or 1, so SkyMask should not be sensitive to threshold $\tau$. We conduct experiments with different thresholds from 0.3 to 0.7, and the results are shown in Table III.
On Fashion-MNIST dataset, it can be seen that SkyMask loses a little accuracy under some attacks when the threshold goes far away from 0.5. Still, it also has a similar accuracy as unattacked FedAvg’s level, with about 1% accuracy loss. SkyMask achieves FNR lower than 1.6%, so there are almost no undetected malicious clients; FPR is lower than 0.9%, so most benign clients are retained.
On CIFAR-10 dataset, especially under LF attack, FPR and FNR increase as the threshold deviates from 0.5. But they are still less than 9%, and the testing accuracy loss is less than 1%. And for HAR, we can see similar results that SkyMask keeps a very low FPR and FNR, achieving the same testing accuracy as unattacked FedAvg. Therefore, the change in threshold does not affect our system’s performance, and SkyMask is not sensitive to the threshold in binarization.
IV-D3 The Impact of Dimension Reduction
The clustering results of SkyMask are influenced by the dimensionality of the features after dimensionality reduction. In Figure 8, we can see that when the number of features after dimensionality reduction rises from 2 to 15, although the testing accuracy does not decrease severely in most cases, there is a noticeable increase in the FPR and FNR, especially under LF attack. The reason is that the framework uses the Gaussian mixture model to cluster the dimension-reduced masks. Since the total number of masks is not large enough, the clustering results may not be satisfying in high-dimension space due to the curse of dimensionality. Therefore, we set the number of features after dimensionality reduction as two empirically.
IV-E SkyMask’s Scalability
To demonstrate the scalability of SkyMask, we conducted experiments utilizing a CNN and ResNet20 global model trained on Fashion-MNIST and CIFAR-10, respectively. For Fashion-MNIST, we assessed the prediction accuracy of the main task, along with the FPR and FNR of SkyMask under various attack scenarios, encompassing 500 and 1000 clients. We have mentioned in §IV-A1 that the local datasets assigned to the clients were disjoint and each client’s dataset should be reasonably sized, so for the CIFAR-10 experiments with fewer data samples, we selected 200 and 500 clients.
Table IV shows that SkyMask can work well even if the number of clients increases to 1000 on Fashion-MNIST. As the number of clients increases, more malicious clients are missed under Min-Sum and Fang-Krum attack, but the FNR is still lower than 1.6%. In addition, the missed malicious local models do not affect the performance of the global model. We can see that the global model of SkyMask achieves the same testing accuracy as the unattacked global model’s obtained by FedAvg.
For the experiments on CIFAR-10, we observed a decrease in the testing accuracy results of the unattacked FedAvg as the total number of clients increased. This can be attributed to the challenges posed by the smaller size of the local datasets and the larger number of clients in the FL training process [23]. However, our SkyMask algorithm demonstrated robust performance even under most attacks. For the LF attack, SkyMask achieved comparable testing accuracy to the unattacked FedAvg, despite a slight increase in the FNR to approximately 9%.
Therefore, SkyMask can work well when the total number of clients increases. Even if there are millions of clients in the FL system, the server can randomly choose a fixed number (e.g., 100 or 500) of clients in each communication round to execute malicious client detection and aggregation. The details we will not describe in this paper.
TABLE IV: The impact of the total number of clients. (The first line in table is the testing accuracy result of FedAvg under no attack.)
FashionMNIST (CNN) Test Acc. FPR FNR
500 1000 500 1000 500 1000
FedAvg 0.89 0.89 / / / /
None 0.89 0.89 / / / /
LF 0.89 0.89 0.40% 0.40% 0.94% 0.62%
Min-Max 0.89 0.89 0.40% 0.30% 0.00% 0.00%
Min-Sum 0.89 0.89 0.40% 1.60% 0.00% 0.00%
Fang-Trim 0.89 0.89 0.00% 0.00% 0.00% 0.00%
Fang-Krum 0.89 0.89 0.60% 1.20% 0.00% 0.00%
Scaling 0.89 / 0.11 0.89 / 0.10 0.00% 0.00% 0.00% 0.00%
CIFAR-10 (ResNet20) Test Acc. FPR FNR
200 500 200 500 200 500
FedAvg 0.75 0.73 / / / /
None 0.75 0.73 / / / /
LF 0.75 0.72 2.94% 6.36% 3.62% 8.36%
Min-Max 0.75 0.74 0.00% 0.00% 0.00% 0.00%
Min-Sum 0.74 0.73 0.00% 0.00% 0.00% 0.00%
Fang-Trim 0.75 0.73 0.00% 0.00% 0.00% 0.00%
Fang-Krum 0.75 0.72 0.00% 0.00% 0.00% 0.00%
Scaling 0.75 / 0.09 0.74 / 0.11 0.00% 0.00% 0.00% 0.00%
V Discussion and Limitations
The current Byzantine attacks are launched with either a single client or multiple clients executing the same attack strategy, which may only account for an increased proportion of the model aggregation stage on the server, but the malicious clients do not collaborate. Increasingly, some researchers find that some specific layers are more vulnerable to Byzantine attack [24]. We also observe that some Byzantine attacks modify parameters near the output layer significantly more than those far from the output layer.
Therefore, each client can be responsible for a different part of model poisoning, using the philosophy of “not putting eggs in the same basket.” The attacker hides poisoning factors in all the uploaded models of the malicious clients it compromises. When most or all malicious model updates bypass the malicious client detection mechanism and get aggregated on the server, it can form a powerful attack on the global model, which is truly a collaborative attack by multiple malicious clients.
The damage to the characteristics of local model updates by this collaborative attack is unknown, and it remains to be verified whether SkyMask can efficiently detect these malicious attacks. For other malicious client detection algorithms, detecting this abnormal pattern in a single model update may be challenging. Theoretically, using our mask mechanism, the collaborative impact of all malicious clients on the global model could be shown after masked aggregation so that the optimization process can change their corresponding masks’ pattern. Actually, Fang attack and AGR-agnostic attack have attempted to attack collaboratively, which weakens the anomaly level of model updates uploaded by each malicious client but influences the performance of the global model when all attacks are combined. Hence, there is a chance for our SkyMask to detect this kind of malicious client.
In the future, we will continue to study this topic in-depth, first verify the effectiveness of this collaborative attack, and deepen the ability of our learnable binary mask to extract the features of model updates and realize the defense capability for this unknown attack. In addition to collaborative attack, the performance of SkyMask under multiple attack types in one communication round is unknown. We should explore how many clusters will the corresponding masks obtained by SkyMask form, and we should also adapt the settings of SkyMask according to this.
Furthermore, SkyMask can be seen as an additional module to the aggregation algorithm. Not only the basic FedAvg algorithm, the SOTA FL algorithms, such as FedProx [25], FedBN [26], or FedAMP [27], can also be equipped with our malicious client detection module. This feature lets the FL system retain the heterogeneity-handling, high convergence rate, or personalization features the original aggregation algorithm may have. It is also possible to enjoy the defensive capability of our malicious client detection strategy, and we will explore the merits of various combinations in future work.
VI Conclusion
We propose a new attack-agnostic robust FL framework called SkyMask to defend against Byzantine attacks. By training parameter-level learnable binary masks on a clean root dataset, SkyMask is the first to implement a fine-grained detection of the poisoned elements of local model updates. Because SkyMask analyzes the mask as a whole, it avoids accidentally removing outliers caused by data heterogeneity on some elements of model updates and precisely removes the malicious clients. Extensive experiments on IID and non-IID data under different attacks prove the effectiveness of SkyMask. Our method shows a solid defensive capability, better than various robust aggregation algorithms and existing malicious client detection methods. Moreover, it also tackles the problem that most clients are malicious and can defend against attacks with a fraction of malicious clients up to $80\%$.

### v2

@@
-SkyMask: Attack-agnostic Robust Federated Learning with Fine-grained Learnable Masks
+
 Abstract
@@
 However, due to the distributed characteristic, FL systems are vulnerable to Byzantine attacks that compromised clients attack the global model by uploading malicious model updates.
-Most existing Byzantine-robust FL systems statistically analyze the weights of whole individual model updates uploaded by clients to defend against Byzantine attacks.
 With the development of layer-level and parameter-level fine-grained attacks, the attacks’ stealthiness and effectiveness have been significantly improved.
-Due to unawareness or overreaction, the existing model-level defense methods degrade the training efficiency and model performance.
-To address this problem, we propose SkyMask, a new attack-agnostic robust FL system that leverages fine-grained learnable masks to identify malicious model updates at the parameter-level.
-Specifically, the FL server applies parameter-level masks to model updates uploaded by clients and trains the masks over a small clean dataset (i.e., root dataset) to learn the subtle difference between benign and malicious model updates in a high-dimension space.
-Our extensive experiments involve different models on three public datasets under state-of-the-art (SOTA) attacks, where the results show that SkyMask achieves up to 10% higher testing accuracy compared with SOTA defense strategies and successfully defends against attacks with malicious clients of a high fraction up to 80%.
-In the meantime, the experimental results demonstrate the scalability of our approach and the weak dependence on the data distribution of the root dataset.
-I Introduction
-Every user click is an oil drop of the digital era that powers up today’s information industry.
-Massive volumes of data generated by users and collected by online services (e.g., social networks, online sales, and streaming platforms) are used to train machine learning (ML) models (e.g., recommendation algorithms) that enhance user experience and boost profits.
-However, with increasingly strict data and privacy policies, such as the GDPR by EU and ATT by Apple, web services tracking and collecting user data for centralized training becomes inappropriate.
-The emergence of federated learning (FL) has addressed such conflicts between ML training and data privacy [1].
-Instead of centralizing training data, FL distributes a model to clients, trains the model locally, and aggregates the local models into a global model without leaking any clients’ local data, which has been applied in many web applications, including online recommendation systems and personal healthcare.
+The existing defense mechanisms solely analyze the model-level statistics of individual model updates uploaded by clients to mitigate Byzantine attacks, which are ineffective against fine-grained attacks due to unawareness or overreaction.
+To address this problem, we propose SkyMask, a new attack-agnostic robust FL system that firstly leverages fine-grained learnable masks to identify malicious model updates at the parameter level.
+Specifically, the FL server freezes and multiplies the model updates uploaded by clients with the parameter-level masks, and trains the masks over a small clean dataset (i.e., root dataset) to learn the subtle difference between benign and malicious model updates in a high-dimension space.
+Our extensive experiments involve different models on three public datasets under state-of-the-art (SOTA) attacks, where the results show that SkyMask achieves up to 14% higher testing accuracy compared with SOTA defense strategies under the same attacks and successfully defends against attacks with malicious clients of a high fraction up to 80%.
+Code is available at https://github.com/KoalaYan/SkyMask.
+1 Introduction
+Federated learning (FL) [22] is a distributed machine learning paradigm that addresses the conflicts between ML training and data privacy.
+Instead of centralizing training data, FL distributes a global model to clients, trains the model locally, and aggregates the local models into a new global model without leaking any clients’ local data, which has been applied in many computer vision applications, including large-scale visual classification [11], object detection [20], medical imaging [10, 18], and many others [35, 19, 17].
 However, due to FL’s distributed design, attackers can easily attack the FL system by compromising client participants and smuggling malicious model updates, known as Byzantine attacks [3, 2, 32, 7, 28, 26].
-Based on whether the compromised clients poison their local data or models, Byzantine attacks can be classified into data poisoning and model poisoning attacks.
-Data poisoning attacks corrupt the global model by injecting contaminated training samples to the local dataset [2, 7].
-Model poisoning attacks [3, 5] directly manipulate local model updates to compromise the global model.
-To tackle these attacks, researchers have explored various defense strategies, most of which leverage coarse-grained model-level statistics, distances, and cosine-similarities to detect outlier model updates (i.e., Multi-Krum [8] and FLTrust [9]) or greedily filter out outlier parameters (i.e., Trim [10] and Bulyan [11]).
-Unfortunately, research on fine-grained attacking methodologies is becoming trendy due to their precise balancing between attacking stealthiness and effectiveness.
-Fine-grained model poisoning attacks [4, 12] leverage the varying sensitivities of different neural network layers and parameters to craft adaptive and stealthy attacks.
-FIBA [13] injects small pixel-wise perturbations in the frequency domain to conduct data poisoning.
+To tackle Byzantine attacks, researchers have explored various defense strategies, most of which leverage coarse-grained model-level statistics to detect outlier model updates [4, 5, 28, 27, 24] or greedily filter out outlier parameters [33, 8].
+Unfortunately, fine-grained attacking methodologies utilize their precise balancing between attacking stealthiness and effectiveness to circumvent existing defense mechanisms, e.g., Fang attack [7] and AGR-agnostic attack [28] leverage the varying sensitivities of different neural network layers and parameters to craft adaptive and stealthy attacks.
 Existing model-level defense strategies either fail to prevent such fine-grained attacks or spill over the benign clients, which extensively sacrifices training efficiency and model quality.
-This paper proposes SkyMask, a new Byzantine-robust FL system that defends against agnostic attacks with fine-grained learnable masks at the parameter-level.
+We conduct a real-world experiment to show the poor performance of the model-level defense [30] under a fine-grained attack [28] (detailed settings provided in supplementary §A.5).
+Existing defense method exhibits 18% testing accuracy degradation.
+We visualize model updates of all clients in a single iteration by the principal component analysis (PCA) as [30].
+As depicted in Fig.
+LABEL:fig:toy_model, the fine-grained attack camouflages the malicious updates within benign ones, underscoring the insufficiency of analyzing local model updates solely at the model level.
+To explore whether binary masks [37, 14] can effectively detect malicious clients, we simultaneously train the binary masks for each model update on the server in the aforementioned toy experiment.
+Fig.
+LABEL:fig:toy_mask illustrates the visualization of trained masks using PCA, where masks applied to malicious model updates distinctly stand out among all masks.
+This highlights the capability of learnable masks to differentiate between benign and malicious model updates, and this method improves the testing accuracy by 18%.
+Based on this observation, we propose SkyMask, a new Byzantine-robust FL system that defends against agnostic attacks with fine-grained learnable masks at the parameter level.
 Specifically, after collecting local models from participant clients, the FL server freezes and multiplies each of the local model parameters with a learnable mask variable.
 Then, all masked local models are trained on a small and clean dataset (i.e., the root dataset [5]) together.
-During the training on the root dataset, the mask variables applied to the frozen local model parameters learn to improve the model accuracy.
-Particularly, the masks applied to malicious models are trained to correct their potential misbehaviors at the parameter-level.
+During the training process on the root dataset, the mask variables, which are applied to the frozen local model parameters, learn to improve the model accuracy.
+Particularly, the masks applied to malicious models are trained to correct their potential misbehaviors at the parameter level.
 Since utilizing fine-grained learnable masks does not enforce customization for specific attacks or refactoring the FL training objective functions, SkyMask gains three superior capabilities beyond existing defense strategies: 1) effective and efficient detection of fine-grained stealthy attacks, 2) attack-agnostic defense against a wide spectrum of attack methods, and 3) strong compatibility and complementarity with existing FL and defense methodology.
-Our main contributions are as follows:
-  • We design fine-grained and computation-efficiency learnable masks that can capture malicious models’ characteristics in a uniform high-dimension space.
-  • We develop SkyMask, a new attack-agnostic Byzantine-robust FL system, which is the first work applying fine-grained learnable masks to detect malicious clients and defend against Byzantine attacks.
-  • We empirically evaluate SkyMask on various benchmarks under six SOTA attacking methods and compare them with existing defense methods, where experimental results show that SkyMask achieves up to 10% higher testing accuracy compared with existing defenses.
-II Preliminaries & Related Work
-II-A Federated Learning
+To sum up, our main contributions are as follows:
+  • We propose fine-grained learnable masks that can capture models’ intrinsic characteristics at the parameter level in a uniform high-dimension space.
+  • We design SkyMask, a new attack-agnostic Byzantine-robust FL system, which applies fine-grained learnable masks to detect malicious clients and defend against Byzantine attacks.
+  • We empirically evaluate our SkyMask on various benchmarks under seven SOTA attacking methods and compare them with existing defense methods, where experimental results show that SkyMask achieves up to 14% higher testing accuracy compared with existing defenses.
+2 Preliminaries & Related Work
+2.1 Federated Learning
 A typical FL system includes $n$ clients and a server.
 Each client $i$ has a local dataset $D_{i}$, $i=1,\dots,n$.
-In each communication round $t$, the server selects a subset of the clients $\mathcal{N}_{t}$ to execute the following three steps: 1) Model distribution: The server distributes the global model $W_{t}$ to the selected clients.
+In each communication round $t$, the server selects a subset of the clients $\mathcal{N}_{t}$ to execute the following steps: 1) Model distribution: The server distributes the global model $W_{t}$ to the selected clients.
 2) Local training: The clients receive global model $W_{t}$ as the initial model $W^{i}_{t+1}$ for local training, and repeat $W_{t+1}^{i}:=W^{i}_{t+1}-\beta\nabla f(W^{i}_{t+1};D_{i})$ for $l$ local iterations, where $f(\cdot;\cdot)$ is the empirical loss function and $\beta$ denotes the local learning rate.
-The corresponding local model update is $\Delta W^{i}_{t+1}=W_{t}-W^{i}_{t+1}$, so for the server, it is equivalent to the local model $W^{i}_{t+1}$.
+The corresponding local model update is $\Delta W^{i}_{t+1}=W_{t}-W^{i}_{t+1}$.
 3) Model aggregation: After the training, each client $i$ uploads its model weights $W_{t+1}^{i}$ to the server for aggregation.
 FedAvg [22] performs weighted model averaging to update the global model: $W_{t+1}=\sum_{i\in\mathcal{N}_{t}}\frac{|D_{i}|}{\sum_{i\in\mathcal{N}_{t}}|D_{i}|}W_{t+1}^{i}$, where we set $|\mathcal{N}_{t}|=n$ for simplicity.
-II-B Threat Model & Byzantine Attacks on FL
-Malicious clients compromised by an attacker participate in the FL process and have access to the knowledge of neural network models, learning rates, and objective functions.
+2.2 Threat Model & Byzantine Attacks on FL
+Malicious clients compromised by the attacker participate in the FL process and have access to the knowledge of neural network models, learning rates, and objective functions.
 Besides, the attacker can fully control these clients’ activities (e.g., modifying the model updates) and has complete access to their local datasets [7, 5, 28].
-Existing Byzantine attacks on FL can be categorized into untargeted, targeted, or backdoor based on the adversary’s goal: 1) Untargeted attacks seek to corrupt the global model and minimize its accuracy on any test input [2, 4, 12, 5].
-For example, Fang [4] adaptively trades off the attacking stealthiness and effectiveness at a fine granularity as an optimization problem.
-2) Targeted attacks aim to minimize the accuracy on the specific test inputs, and maintain high accuracy on the rest [3, 14].
-3) Backdoor attacks reduce the utility of the global model on test inputs that contain a specific signal called trigger [15, 6].
-II-C Byzantine-robust FL Algorithms
+Existing Byzantine attacks on FL can be categorized into untargeted and targeted based on the adversary’s goal: 1) Untargeted attacks seek to corrupt the global model and minimize its accuracy on any test input [30, 7, 28, 2].
+Specifically, fine-grained attacks, represented by Fang attack [7] and AGR-agnostic attack [28], adaptively trade off the attacking stealthiness and effectiveness at a fine granularity as an optimization problem, which preserves the effects of poisoning with as few model-level outliers as possible.
+2) Targeted attacks aim to reduce the utility of the global model on attacker-specified tasks [3, 29, 1, 32, 15, 21, 25, 34].
+2.3 Byzantine-robust FL Algorithms
 We classify existing representative Byzantine robust algorithms into two categories based on their strategies:
-II-C1 Model-level defense strategies
-FLDetector [16] determines whether a client is malicious by performing the consistency analysis on its model updates.
-Specifically, FLDetector predicts a client’s model update for each iteration based on historical records and compares it with the received model update.
-If they are inconsistent in multiple iterations, the client is probably malicious.
-FLTrust [9] trains a root model with a small clean root dataset at the server in each round, and measures the cosine similarity between a local model and the root model.
-The similarity metric dictates the weight in global model aggregation.
-Tolpegin [2] detects malicious participants by analyzing the characteristics of model updates.
-It standardizes all model updates, uses the principal component analysis (PCA) for dimension reduction, and identifies malicious clients via clustering algorithms.
-Krum [8] assumes the number of malicious clients is known as $n_{m}$.
-The server computes a score $s^{i}_{t+1}$ for each model update from the sum of squares of the distance between this model update and the $n-n_{m}-2$ nearest ones.
-Then, the one with the smallest score is selected as the global model update for the next round.
-Multi-Krum (MKrum) [8], a Krum variant, selects $c$ model updates with the lowest scores, and then aggregates them for a new global model.
-II-C2 Greedy parameter-level filtering strategies
-Trimmed-Mean (Trim) and Median [10] are coordinate-wise aggregation rules assuming the number of malicious clients $n_{m}$ is known, remove the parameters of the smallest and largest $n_{m}$ values, and average the remaining.
-Median retains the corresponding coordinate’s median value as the new global model’s value.
-Bulyan [11] also relies on a known $n_{m}$ and requires $n\geq 4n_{m}+3$ to ensure its robustness.
-It first iteratively applies MKrum to select $\theta$ local models, where $\theta<n-2n_{m}$, and uses a variant of Trim to aggregate the $\theta$ selected local models.
-However, existing robust FL strategies can hardly defend against those fine-grained attacks without sacrificing training efficiency and model quality by utilizing coarse-grained similarity checking or greedy parameter-wise filtering, which motivates us to explore malicious client detection with fine-grained learnable masks.
-II-D Motivating Learnable Masks
-(a) The PCA of model updates
-(b) The PCA of masks
-Figure 1: Visualizing model updates and masks with PCA.
-Binary masks have been widely used in deep learning and demonstrated the capability to capture neural network models’ intrinsic characteristics at the parameter-level.
-For example, Zhou et al[17] prune deep neural network models with binarized masks to improve training efficiency.
-FedMask [18] utilizes learnable binary masks to enable each device to learn a personalized and structured sparse neural network model that runs efficiently on devices.
-We use a realistic experiment to show how learnable masks can be applied to detect malicious clients.
-In a swarm of 100 FL clients, we compromise 20 as malicious clients to apply Min-Sum attack.
-We train a four-layer CNN model on Fashion-MNIST dataset with the non-IID setting described in §IV-A1.
-We visualize the CNN model updates of all clients in one iteration using PCA in Figure 1(a), which shows how the fine-grained Min-Sum attack hides the malicious model updates among the benign updates.
-Hence, analyzing local model updates with PCA is insufficient to distinguish benign and malicious clients.
-To detect such stealthy attacks, we apply fine-grained learnable masks that capture the parameter-level characteristics of local models.
-We aggregate all masked local models and train the masks together on a small root dataset, where the learnable masks learn to enhance the benign models and correct the malicious models (details in §III-B).
-Figure 1(b) shows the visualization of trained masks using PCA, where the masks applied to the malicious model updates are easily distinguished among all masks.
-This phenomenon reveals the capability of learnable masks that captures the difference between benign and malicious model updates.
-II-E Root Dataset
-Considering that the server has no root of trust to decide whether a model update is malicious or not, FLTrust [9] was the first to introduce the concept of root dataset.
-It can take advantage of the root dataset even if the distribution diverges from the overall data distribution or the size is less than one hundred.
-Therefore, it is easy for the server and FL manager to manually collect or generate a small root dataset.
-The rationality for using root datasets has been verified in several other approaches [19, 20].
-III SkyMask’s Design
-III-A Goals
+Model-level defense strategies.
+Some defense methods take model-level distance as a basis for determining whether a client is malicious, e.g., Euclidean distance in Krum [4], or cosine distance in DeepSight [27] and FLAME [24].
+FLTrust [5] trains a root model with a small root dataset, and takes the cosine similarity between a local model and the root model as the weight in aggregation.
+Tolpegin defense [30] identifies malicious model updates by employing PCA dimension reduction and clustering.
+FLDetector [36] predicts a client’s model update based on historical data and compares it with the received update.
+It employs consistency analysis on model updates to identify malicious clients.
+Greedy parameter-level filtering strategies.
+Trimmed-Mean (Trim) [33] is a coordinate-wise aggregation rule assuming the number of malicious clients $n_{m}$ is known, removes the parameters of the smallest and largest $n_{m}$ values, and averages the remaining.
+However, existing robust FL strategies can hardly defend against those fine-grained attacks without sacrificing training efficiency and model quality by utilizing coarse-grained detection or greedy parameter-wise filtering, which motivates us to explore malicious client detection with fine-grained learnable masks.
+2.4 Root Dataset
+Considering that the server has no root of trust to decide whether a model update is malicious or not, FLTrust [5] is the first to introduce the concept of root dataset.
+It can take advantage of the root dataset even if the distribution diverges from the overall data distribution or the size is less than one hundred, which is easy for the server and FL manager to manually collect or generate a small root dataset.
+The rationale for utilizing root datasets has been validated in numerous other studies [9, 23, 31, 16].
+We evaluate the impact of the root dataset’s data distribution in supplementary §C.1.
+3 Methodology
 We aim to design a server-based robust FL system.
-The server has a small representative root dataset following a previous work [9].
-Moreover, this root dataset can be imbalanced without a major bias (analyzed in §IV-D1).
-We allow reasonable additional computation costs and extra storage space to defend against Byzantine attacks.
-We aim to detect all the malicious model updates and filter them out at each communication round to eliminate the impact of attacks.
-We should also keep as many benign clients as possible in the filtered FL client set to train a final global model with sufficient local data for satisfactory accuracy.
-For the proposed Byzantine-robust FL system, it should promise the following three features:
-  • Robustness.
-The system should preserve the global model accuracy when attacked by malicious clients.
-In particular, we should maintain a high malicious client detection accuracy and a low misidentification rate of benign ones (i.e., false positive rate).
+The server has a small representative root dataset following previous work [5].
+Moreover, this root dataset can be imbalanced.
+For the proposed Byzantine-robust FL system, it should promise the following three features: (1) Robustness.
+The system should preserve the global model accuracy under attacks.
+In particular, we should maintain a high malicious client detection accuracy and a low misidentification rate of benign ones to eliminate the impact of attacks.
+(2) Generality.
 The system should also stay effective on different datasets and model structures under different attacks.
-  • Efficiency.
-The system should not introduce extra computation and communications overhead to the clients, which are often resource-constrained devices in FL.
-Besides, when performing malicious client detection and defense on the server, the system should only incur a proper computation complexity.
-  • Scalability.
-The system should still support a large scale of clients and work under a high fraction of malicious clients (e.g., 80%) since FL systems usually involve a large number of client devices.
-III-B SkyMask: Overview
+(3) Efficiency.
+The defense system is designed on the server instead of resource-constrained client.
+Besides, when performing malicious client detection and defense on the server, we allow reasonable additional computation costs and extra storage space.
+3.1 Overview
+Figure 4: SkyMask’s workflow.
 When initializing the FL process, the server creates an initial global model and builds a root dataset as [5].
-Figure 2 presents SkyMask’s workflow, where each communication round has six steps: Step 1, the server distributes the global model parameters to each client.
-Step 2, the clients load the parameters, train the local models and send back these model updates.
-Step 3, the server freezes received local models and applies learnable masks to each local model update.
-Step 4, all the masked local models are trained together on the root dataset to converge.
-Step 5, the server detects and removes the malicious clients by clustering all trained masks.
-Step 6, the server aggregates the remaining model updates into a new global model.
-The key component of SkyMask is the detection mechanism based on learnable masks.
+Fig.
+4 presents SkyMask’s workflow, where each communication round has six steps: Step 1, the server distributes the global model parameters to each client; Step 2, the clients load the parameters, train the local models, and send back these model updates; Step 3, the server freezes and multiplies each local model parameter with a learnable mask variable; Step 4, all the masked local models are trained together on the root dataset to converge; Step 5, the server detects and removes the malicious clients by clustering all trained masks; Step 6, the server aggregates the remaining model updates into a new global model.
+The key component of SkyMask is the learnable mask-driven detection.
 Specifically, SkyMask constructs a same-size fine-grained mask for each local model update and trains these masks by optimizing the aggregation result of masked local models until they converge.
@@
 After removing all the detected malicious model updates, SkyMask calculates the average of remaining model updates and uses it to update the global model.
-III-C The Complete SkyMask Algorithm
+3.2 SkyMask Algorithm
 Most existing defense strategies are based on the local model updates’ coarse-grained statistics or greedily filtering out outlier parameters.
@@
 Our defense strategy has two stages: 1) mask initialization and training stage; 2) mask clustering and classification stage.
-Figure 2: SkyMask’s workflow.
-III-C1 Mask initialization and training
-In the first stage, the server assigns a learnable mask $m_{i}$ of the same size as the model for each client $i$ and initializes them all with $1$s.
+3.2.1 Mask initialization and training.
+In the first stage, the server freezes all the model updates and assigns a learnable mask $m_{i}$ of the same size as the model for each client $i$ and initializes them all with $1$s.
 The model aggregated from masked local models $\tilde{W}_{t+1}$ (called aggregated masked model) is computed by averaging the masked local models $m_{i}\odot W^{i}_{t+1}$.
@@
 At the beginning of iteration $t+1$, $\tilde{W}_{t+1}$ is calculated as:
-$W^{i}_{t+1}=W_{t}-\Delta W^{i}_{t+1},$
 $\tilde{W}_{t+1}=\sum^{n}_{i=1}\frac{1}{n}m_{i}\odot W^{i}_{t+1}.$ (1)
@@
 Then, the update process back-propagates gradients with the help of $\tilde{W}_{t+1}$ to masks and applies a gradient descent algorithm to the masks.
-Intuitively, we know that if we fix all masks as zeros except for one mask to be trained, there is also a solution for this optimization problem.
-If $\tilde{W}_{t+1}$ converges to the fixed optimal model $\hat{\tilde{W}}_{t+1}$, the problem is equivalent to solving a multivariate primary equation, which has infinite solutions.
-Moreover, suppose the mask value is not limited.
-In that case, some masks have very large values, and others are so small for the same dimension that these masks can not effectively characterize whether the local model is benign.
-So we cannot detect malicious clients with the help of this type of mask, and we should redesign the masks to show the contribution of each local model update in the optimized model.
-Binary can directly determine whether a parameter of model update participates in the global model, and binary masks have shown effectiveness in model pruning.
-So we use binary masks to extract the parameter-level characteristics of local models.
-If the malicious model parameters attempt to corrupt the aggregated model, the optimization process prefers to reject them and sets the corresponding parameters of the mask as 00.
-III-C2 Approximation of binary masks
-Binary parameters cannot be easily optimized by a gradient descent algorithm.
-Thus, we approximate the binary masks in training by setting the mask with real values and applying a sigmoid function $\sigma(\cdot)$ instead of a hard threshold to reduce the enormous gradient variance.
+If there is no constraint on the value range of mask variables during the mask training, some masks can have very large values, while others can be so small for the same dimension when these masks converge.
+Huge differences in magnitude between masks can affect the detection of malicious clients.
+Therefore, we redesign the masks to show the contribution of each local model update in the optimized model.
+We use binary masks to extract the parameter-level characteristics of local models because binary can directly determine whether a parameter of model update participates in the global model.
+Since binary parameters are not derivable, we approximate the binary masks in training by setting the mask with real values and applying a sigmoid function $\sigma(\cdot)$ instead of a hard threshold to reduce the enormous gradient variance [14].
 Furthermore, it is easy for the gradient to be back-propagated to the real-value masks.
-The approximation of a binary mask is calculated as $\tilde{m}_{i}=\sigma(m_{i})$.
-Since the mask’s limitation is changed, the original Equation 1 is unsuitable.
-To adapt for the approximated binary masks, $\tilde{W}_{t+1}$ is re-written as:
+The approximation of a binary mask is calculated as $\tilde{m}_{i}=\sigma(m_{i})$, which is ranged $(0,1)$.
+Since the mask’s limitation is changed, the original Equation 1 is unsuitable. To adapt for the approximated binary masks, $\tilde{W}_{t+1}$ is re-written as:
 $\tilde{W}_{t+1}=\frac{\sum^{n}_{i=1}\tilde{m}_{i}\odot W^{i}_{t+1}}{\sum^{n}_{i=1}\tilde{m}_{i}}.$ (2)
@@
 Thus, the 0-1 pattern in the mask can represent whether the corresponding local model is malicious.
-The mask training task is not different from the main task in FL except for the variables. By using the stochastic gradient descent algorithm (SGD) to optimize these masks, the objective function is:
+The mask training task is not different from the main task in FL except for the variables. The objective function is:
 $f(\tilde{W}_{t+1},D_{r})=\sum\limits_{(x,y)\in D_{r}}L(\text{output}(x,\tilde{W}_{t+1}),y).$ (3)
-During training the masks, masks are updated in each mask training iteration as follows:
+The masks are updated in each mask training iteration as follows:
 $m_{i}:=m_{i}-\gamma\cdot\nabla_{m_{i}}f(\tilde{W}_{t+1},D_{r}),$ (4)
 where $\gamma$ represents the mask learning rate.
-This process will not stop until these masks converge. After the real-value masks converge, the final results of binary mask $\hat{m}_{i}$ are computed as Equation 5.
+This process will not stop until these masks converge.
+After the real-value masks converge, the final results of binary mask $\hat{m}_{i}$ are computed as:
 $\hat{m_{i}}[k]=\begin{cases}1,&\tilde{m}_{i}[k]>\tau\\ 0,&\tilde{m}_{i}[k]\leq\tau\end{cases},$ (5)
 where $\hat{m}_{i}[k]$ represents the parameter located in $k$th dimension of binary mask $\hat{m}_{i}$ of client $i$ and $\tau$ is a threshold.
-III-C3 Mask clustering and classification
-These binary masks can sense the fine-grained poisoning attack and represent the parameter-level characteristics of each client’s model update.
-We use the PCA algorithm to reduce the dimensionality to $d$.
-Since the maximum dimensionality that the masks can be reduced to is limited by the number of masks $n$ and the dimensionality of masks $V$, we set $d=2$, which is the minimum possible number of clients in the FL system.
+3.2.2 Mask clustering and classification.
+These binary masks can sense potential attacks at a fine granularity and represent the parameter-level characteristics of each client’s model update.
 Then, we apply the Gaussian mixture model for clustering and classification.
 If there are no attacks, the clustering result is only one cluster, and the server aggregates all the benign local models.
-If the fraction of malicious clients is lower than 50%, the system can treat the larger cluster in two clusters as the benign client set.
-To cope with more general situations, SkyMask introduces a trusted root model $W^{r}_{t+1}$, which is initialized by $W_{t}$ and “locally” trained $l$ iterations on the root dataset as other clients do:
+To determine which cluster is benign, SkyMask introduces a trusted root model $W^{r}_{t+1}$, initialized by $W_{t}$ and “locally” trained $l$ iterations on the root dataset as other clients:
 $W^{r}_{t+1}:=W^{r}_{t+1}-\beta\nabla f(W_{t},D_{r}).$ (6)
-Then the server lets it join in the model set before the mask training stage, assigns one mask for it and trains this mask together with other masks.
+The server lets it join in the model set before the mask training stage, assigns one mask for it and trains this mask together with other masks.
 When all the masks converge, the clients represented by the masks in the same cluster as the mask corresponding to the trusted root model are considered benign.
-This way, SkyMask can handle various malicious fractions from zero to very high.
-In order not to interfere with the global model convergence, we should emphasize that the trusted root model is only used for malicious model detection and will not be included in the benign model aggregation phase.
-SkyMask plays $R$ rounds for server-clients communication, and each round contains six steps, as §III-B describes.
+In this way, SkyMask can handle various malicious fractions from zero to very high.
+In order not to interfere with the global model convergence, we should emphasize that the trusted root model is only used for malicious model detection and is not included in the benign model aggregation phase.
+SkyMask plays $R$ rounds for server-clients communication, and each round contains six steps, as §3.1 describes.
 The critical point of SkyMask is our fine-grained malicious client detection strategy mentioned above, which can provide the remaining benign client set $U_{b}$ for aggregating a new global model.
-The global model for the next communication round $W_{t+1}$ is calculated as:
+Given the aggregation algorithm specified as FedAvg, the global model for the next communication round $W_{t+1}$ is calculated as:
 $W_{t+1}=W_{t}-\alpha\cdot\sum\limits_{i\in U_{b}}\frac{|D_{i}|}{\sum_{i\in U_{b}}|D_{i}|}\Delta W^{i}_{t+1},$ (7)
 where $\alpha$ denotes the global learning rate.
-It should be noted that SkyMask is more concerned with malicious client detection and can be added as an additional module to the aggregation algorithm.
-Therefore, it can even enhance the existing Byzantine-robust aggregation algorithms.
 If it completes the malicious detection task, removing only the malicious model updates and leaving the benign model updates untamed, its convergence will be consistent with the original aggregation algorithm.
-Figure 3: Visualizing the training process of masks applied to malicious and benign models, respectively.
-III-D Why SkyMask Works
-We conduct an example experiment to show how SkyMask works.
-We choose a four-layer CNN model and Fashion-MNIST dataset and set 20% of 100 clients as malicious.
-Under Fang-Trim attack, we take the mask training process in one server-client communication round.
-We calculate the proportion of “1”s in each part of the binary mask, and a smaller value means fewer parameters of this client are selected for the aggregated masked model.
+SkyMask serves as an efficient method for identifying malicious clients within an FL framework.
+Its modular design allows for integration as an adjunct component into diverse aggregation algorithms, ensuring compatibility across a spectrum of existing aggregation techniques.
+3.3 Why SkyMask Works
+Figure 5: The visualization of the training process of masks applied to malicious and benign models.
+We choose a four-layer CNN model and Fashion-MNIST dataset, and set 20 of 100 clients as malicious.
+Under Fang-Trim attack, we take the mask training process in one communication round.
+We calculate the proportion of “1”s in each part of the binary mask.
+A smaller value means a fewer parameters of this client are selected for the aggregated masked model.
 For simplicity, here only shows 20 masks for malicious and 20 for benign clients in each mask training iteration.
-As Figure 3 shown, during the mask training process, the binary parameters in the “conv1.weight”, “conv2.weight” and “fc1.weight” parts of malicious model updates are set to zero more than those of benign updates.
-It is related to the fact that the attack prefers to attack the parameters in specific parts.
-The masks gradually reduce the poisoning effect of these toxic parameters during the training so that we can classify malicious clients with the help of these masks.
-Therefore, SkyMask can perceive the poisoning attacks at the parameter-level in the model updates without knowing the details of the attack, which makes it attack-agnostic.
-III-E Complexity
+As shown in Fig.
+5, during the mask training process, the binary parameters in the “conv1.weight”, “conv2.weight”, and “fc1.weight” parts of malicious model updates are set to zeros, which are more often than those of benign updates.
+It is related to the fact that the attack prefers to poison the parameters in specific parts.
+The masks gradually reduce the poisoning effect of the toxic parameters during the training so that we can classify malicious clients with the help of the masks.
+Therefore, SkyMask can perceive the poisoning attacks at the parameter level in the model updates without knowing the details of the attack, which makes it attack-agnostic.
+3.4 SkyMask’s Complexity
 Let $O(T)$ denote the computational complexity of one local training iteration, and the model parameters’ dimensionality is $V$.
-To update the masks once, the server calculates the gradient of each mask from the gradient of the aggregated masked model $\tilde{W}_{t+1}$ by backpropagation, and its computational complexity is $O(nV)$.
-Hence, the total computational complexity is $O((T+nV)t_{m})+O(nV)=O(Tt_{m}+nVt_{m})$, where $t_{m}$ is the number of mask training convergence iterations.
+The computational complexities of existing defense methods are as follows: $O(nV)$ for FLTrust, $O(nV\log n)$ for Trim, $O(n^{2}V)$ for Krum, $O(T+nV)$ for DeepSight, and $O(nV)$ for FLAME.
+For SkyMask, to update the masks once, the server computes the gradient of each mask from the gradient of the aggregated masked model $\tilde{W}_{t+1}$ by backpropagation, with a computational complexity of $O(nV)$.
+Hence, the total computational complexity of SkyMask is $O((T+nV)t_{m})+O(nV)=O(Tt_{m}+nVt_{m})$, where $t_{m}$ is the number of mask training convergence iterations.
 Since there is no data dependency between any mask parameters, we can use parallel training [6] to accelerate the mask optimization process so that the time overhead can be further reduced, trading computation power for time.
-FLTrust only needs to calculate the similarity of local models and its baseline model, whose computational complexity is $O(nV)$.
-Trim sorts each parameter of local models and choose the middle part of the values, so its computational complexity is $O(nV\log n)$.
-Krum calculates the distance score of each local model, which is the sum of the distance between one local model and its $n-n_{m}-2$ nearest local models, so its computational complexity is $O(n^{2}V)$.
-The above analysis shows that the additional cost of the mask learning mechanism over other defense methods is equivalent to some iterations of local training, which are acceptable compared to the computational power of the server.
-The space complexity of existing FL defense algorithms is either $O(n+nV)$ or $O(n^{2}+nV)$.
-Since we assign a mask for each client, the additional space complexity is also proportional to the number of clients $n$.
-It is proportional to the space complexity of the global model parameters because the masks have the same dimensionality as the global model parameters.
-Considering that the server stores $n$ local models’ parameters in the current round, the total space cost is about twice the others, and its space complexity is $O(nV+nV)=O(nV)$, which is reasonable.
-IV Evaluation
-We implement a prototype of SkyMask based on PyTorch and evaluate our algorithm with existing Byzantine-robust FL aggregation algorithms and two malicious client detection algorithms.
-To verify the generalization of our method, we conduct experiments using several popular datasets and pair them with different models, respectively.
-We evaluate our approach with six popular attacks, explore the impact of the fraction of malicious clients, and verify its scalability.
-In sensitivity analysis, we discuss the influence of binarization threshold value, the number of features after dimensionality reduction, and the root dataset’s data distribution on our detection method’s performance.
-We repeat each experiment three times and average the results.
-All experiments run on a cloud server with eight NVIDIA RTX 2080 GPUs.
-On Fashion-MNIST and HAR datasets, SkyMask improves the testing accuracy by 1–3% compared with the best result of all five Byzantine-robust algorithms under the same attacks.
-More significantly, our method gains a 10% improvement in testing accuracy on CIFAR-10 dataset under Min-Max attack.
-IV-A Experiment Settings
-IV-A1 IID and non-IID Datasets
-Following previous work [4, 9], in different datasets for classification tasks, if there are $C$ classes, we construct $C$ groups.
-The probability of a training sample with label $k$ assigned to group $k$ is $q>0$ and to other group is $\frac{1-q}{C-1}$.
-Each client randomly chooses one group and samples a certain number of samples uniformly from this group for its local dataset.
-The probability $q$ controls the degree of non-IID.
-If $q=\frac{1}{C}$, the dataset for each client is IID.
-For the root dataset, we randomly sample a certain number of samples from the training dataset with a bias probability $p$ whose meaning is similar to $q$ and then split the rest for each client in the way mentioned above.
-We use Fashion-MNIST, CIFAR-10, and HAR dataset [22] as the evaluation datasets, which have been widely used in prior FL studies.
-The Fashion-MNIST dataset consists of a 60K image samples training set and a 10K samples test set, where each image is associated with a label from 10 classes.
-The CIFAR-10 dataset composes 50K training images and 10K test images in 10 classes.
-They are colorful with three channels, contain many features noise, and have varying object scales.
-For both Fashion-MNIST and CIFAR-10, we use $q=0.5$ to conduct the non-IID experiments as [4].
-By setting $q=0.1$, we also conduct the IID experiments as [12].
-HAR dataset is built by recording 30 participants’ daily living with sensors, consisting of 7,352 pieces of 561-feature vectors with time and frequency domain variables.
-The main task is to classify the activity into one of six human activities.
-Since they are collected naturally by 30 devices, it is unnecessary to divide the dataset manually.
-We randomly assign 70% of each device’s dataset as the training dataset and 30% to be the testing data.
-IV-A2 Models
+The above analysis shows that the additional cost of the mask learning mechanism over other defense methods is equivalent to the cost of conducting several iterations of local training, which is acceptable compared to the computational power of the server.
+The space complexity of existing defense methods are: $O(nV)$ for Trim, $O(n+nV)$ for FLTrust, $O(n^{2}+nV)$ for Krum, DeepSight, and FLAME.
+For SkyMask, the additional space complexity is proportional to the number of clients $n$ and the space complexity of the global model parameters because SkyMask assigns a mask for each client and the masks have the same dimensionality as the global model.
+Considering that the server stores $n$ local models’ parameters in the current round, the total space cost is about twice the others, and its space complexity is $O(nV)$, which is reasonable.
+4 Experiments
+We evaluate our method with existing Byzantine-robust FL aggregation algorithms and two malicious client detection algorithms.
+To verify the generalization of our method, we conduct experiments using various popular datasets paired with different models.
+We evaluate our method with seven popular attacks, explore the impact of the fraction of malicious clients, and verify its scalability.
+In sensitivity analysis, we discuss the influence of binarization threshold, the number of features after dimensionality reduction, and root dataset’s data distribution on our method’s performance (details in supplementary §C), respectively.
+4.1 Experiment Settings
+4.1.1 Datasets and models.
+We evaluate SkyMask on three widely-used datasets: Fashion-MNIST, CIFAR-10, and CIFAR-100 [13].
+Following previous methods [7, 5], we construct non-IID local dataset with bias probability $p=0.5$ for each client and the root dataset for the server.
 We construct a CNN with two convolutional layers and two linear layers as the global model for Fashion-MNIST dataset.
-For CIFAR-10 dataset, to verify SkyMask’s effectiveness on large models in FL, we use a more complex model, ResNet20.
-We select a linear regression classifier for HAR dataset.
-IV-A3 Baseline attacks
-The experiments cover untargeted, targeted, and backdoor attacks, including the following baselines:
-  • Fang attack [4]: Fang attack maximizes the deviation of the global model towards the opposite direction of the unattacked global model update.
-By using Trim and Krum as the aggregation algorithm in the attack, we obtain two different attacks: Fang-Trim and Fang-Krum.
-  • Label-flipped (LF) [2]: Following the same setting as Fang [4], we flip label $k$ to $C-k-1$ for each sample on malicious clients, where $k\in\{0,\dots,C-1\}$ and $C$ denotes the total number of dataset classes.
-  • AGR-agnostic [12], i.e., Aggregation algorithm agnostic attack, aims to limit the range of malicious updates and maximize the perturbation.
-It has two variants: Min-Max attack restricts the range based on the maximum value of the L-$2$ distance between any pair of benign model updates.
-Min-Sum attack is based on the minimum value of the sum of squares of the distances between any benign model update and all other benign model updates.
-  • Scaling attack [15]: Scaling attack is a targeted backdoor attack.
-It poisons the local dataset with trigger-embedded samples and scales up the model updates before sending them to the server.
-We use the same pattern trigger in LIE [5] and follow the same data augmentation scheme as FLTrust [9].
-IV-A4 Baseline defenses
-In §IV-B, to show the defense effectiveness, we compare SkyMask with five Byzantine-robust FL aggregation algorithms: FLTrust [9], Trim [10], Krum, MKrum [8] and Bulyan [11].
-In §IV-C, to reveal the significance of learnable masks in malicious client detection, we compare SkyMask with two malicious client detection algorithms: Tolpegin defense [2] and FLDetector [16].
-TABLE I: FL testing accuracy under different attacks and attack success rates of Scaling attack. The experimental results of Scaling attack is in the form of “testing accuracy / attack success rate”.
-Dataset (Model) Attack FedAvg FLTrust Trim Krum MKrum Bulyan SkyMask-NR SkyMask
-Fashion -MNIST (CNN) non-IID None 0.89 0.89 0.88 0.83 0.88 0.87 0.89 0.89
+For CIFAR-10 and CIFAR-100 datasets, to verify SkyMask’s effectiveness on large models in FL, we use a more complex model, ResNet20.
+4.1.2 Baseline attacks and defenses.
+The experiments cover 1) untargeted attacks: Label-flipped (LF) [30] and four fine-grained attacks, including Fang attacks [7] (Fang-Trim and Fang-Krum attacks) and two types of the AGR-agnostic attack [28] (Min-Max and Min-Sum attacks); 2) targeted attack: Scaling attack [1] and DBA [32].
+In §4.2, to show the defense effectiveness, we compare SkyMask with five Byzantine-robust FL aggregation algorithms: FLTrust [5], Trim [33], Krum [4], DeepSight [27] and FLAME [24].
+In §4.3, to reveal the significance of learnable masks in malicious client detection, we compare SkyMask with two malicious client detection algorithms: Tolpegin defense [30] and FLDetector [36].
+4.1.3 FL parameter settings.
+Following the previous study [5], we use 100 clients in all experiments and apply multiple-client attacks.
+The default fraction of malicious clients is set to 20%.
+In our experiments, the root dataset has 100 samples and no intersection with the training and test datasets, which is similar to FLTrust [5].
+Other hyperparameters are detailed in supplementary §A.4.
+4.1.4 SkyMask variants.
+Intuitively, if the fraction of malicious clients is less than 50% in the real world, we can choose the larger one in two clusters of masks as benign.
+This way, we can further reduce the computational pressure on the server.
+To verify the effectiveness with or without the trusted root model, we consider the following two schemes: 1) SkyMask-NR (no trusted root model).
+The FL system does not generate a trusted root model, and after all the masks are converged, the system chooses the larger one in two clusters of the masks as benign for aggregation; 2) SkyMask.
+The FL system uses the defense strategy fully following the default steps described in §3.2.
+4.1.5 Metrics.
+The testing accuracy is obtained from the prediction accuracy of the global model on the test dataset.
+The attack success rate under targeted attack is the fraction of the trigger-embedded samples in the test dataset that are identified as the target label.
+To show the malicious client detection ability, we record false positive rate (FPR) and false negative rate (FNR) in the malicious client detection stage.
+FPR indicates the number of benign clients that are misidentified as malicious clients divided by the total number of benign clients, and FNR indicates the number of malicious clients that are misidentified as benign clients divided by the total number of malicious clients.
+The Mean FPR/FNR is the average of FPR/FNRs in each communication round.
+4.2 The Defense Effectiveness of SkyMask
+We first certificate the defensive capability under a low fraction of attacks.
+Then, there is also possibly no attacker in the FL system, so the defense methods should not affect the global model’s performance.
+Eventually, we consider an extreme condition that the fraction of malicious clients is high.
+Table 1: FL testing accuracy under different attacks and attack success rates of targeted attack.
+The experimental results of targeted attacks are in the form of “testing accuracy/attack success rate.”
+Dataset (Model) Attack FedAvg FLTrust Trim Krum DeepSight FLAME SkyMask-NR SkyMask
+Fashion -MNIST (CNN) None 0.89 0.89 0.88 0.83 0.88 0.87 0.89 0.89
 LF 0.84 0.86 0.84 0.83 0.89 0.87 0.89 0.89
@@
 Scaling 0.80/0.21 0.89/0.08 0.87/0.13 0.82/0.09 0.88/0.10 0.87/0.10 0.89/0.10 0.89/0.10
-Fashion -MNIST (CNN) IID None 0.90 0.89 0.89 0.86 0.89 0.87 0.89 0.90
-LF 0.81 0.88 0.84 0.86 0.89 0.87 0.89 0.90
-Min-Max 0.63 0.89 0.76 0.86 0.59 0.72 0.90 0.90
-Min-Sum 0.85 0.89 0.78 0.55 0.84 0.68 0.90 0.90
-Fang-Trim 0.53 0.89 0.74 0.86 0.84 0.86 0.90 0.90
-Fang-Krum 0.88 0.89 0.86 0.26 0.71 0.78 0.90 0.90
-Scaling 0.83 / 0.21 0.89 / 0.10 0.88 / 0.14 0.85 / 0.11 0.89 / 0.11 0.87 / 0.11 0.89 / 0.10 0.90 / 0.10
-CIFAR-10 (ResNet20) non-IID None 0.77 0.75 0.77 0.54 0.72 0.72 0.76 0.76
+DBA 0.86/0.51 0.88/0.18 0.88/0.31 0.83/0.11 0.89/0.16 0.89/0.11 0.89/0.11 0.89/0.10
+CIFAR-10 (ResNet20) None 0.77 0.75 0.77 0.54 0.72 0.72 0.76 0.76
 LF 0.71 0.74 0.71 0.55 0.69 0.68 0.75 0.76
@@
 Scaling 0.10/1.00 0.74/0.13 0.72/0.51 0.53/0.08 0.73/0.10 0.73/0.10 0.77/0.09 0.77/0.10
-CIFAR-10 (ResNet20) IID None 0.79 0.78 0.78 0.63 0.73 0.73 0.79 0.79
-LF 0.76 0.76 0.74 0.64 0.72 0.73 0.77 0.78
-Min-Max 0.68 0.61 0.58 0.56 0.65 0.69 0.79 0.79
-Min-Sum 0.73 0.75 0.62 0.30 0.66 0.70 0.79 0.79
-Fang-Trim 0.10 0.70 0.38 0.62 0.54 0.69 0.78 0.78
-Fang-Krum 0.68 0.77 0.60 0.42 0.34 0.34 0.79 0.79
-Scaling 0.10 / 1.00 0.76 / 0.12 0.75 / 0.67 0.62 / 0.10 0.74 / 0.10 0.74 / 0.10 0.78 / 0.11 0.79 / 0.10
-HAR (LR) None 0.96 0.96 0.96 0.81 0.95 0.94 0.96 0.96
-LF 0.88 0.95 0.81 0.86 0.79 0.74 0.96 0.96
-Min-Max 0.89 0.95 0.93 0.76 0.83 0.66 0.96 0.96
-Min-Sum 0.96 0.95 0.94 0.30 0.86 0.83 0.96 0.96
-Fang-Trim 0.23 0.96 0.80 0.77 0.65 0.82 0.96 0.96
-Fang-Krum 0.96 0.96 0.93 0.15 0.75 0.71 0.96 0.96
-Scaling 0.72 / 0.20 0.96 / 0.20 0.92 / 0.16 0.84 / 0.19 0.96 / 0.20 0.94 / 0.19 0.96 / 0.19 0.96 / 0.20
-IV-A5 FL parameter settings
-We use 100 clients in Fashion-MNIST and CIFAR-10 experiments, and in HAR, we use 30 clients, following previous study [9].
-To boost the effectiveness of attacks, we apply multiple-client attacks instead of single-client attacks, and the default fraction of malicious clients is set to 20%.
-The default binarization threshold $\tau$ is $0.5$, and the root dataset bias probability $p=0.1$.
-We set both global model learning rate $\alpha$ and local model learning rate $\beta$ to $0.5$ in all experiments.
-For mask learning rate $\gamma$, we set $1\times 10^{7}$ in Fashion-MNIST and HAR datasets but $1\times 10^{8}$ in CIFAR-10 dataset.
-This is because the order of magnitude of the gradient is $10^{-8}-10^{-7}$, which declines along the model’s dimensionality increases, so we should set it larger to help masks converge.
-In CIFAR-10, we allow the system to execute server-client communication for $500$ rounds and let clients do $l=5$ local training iterations per round.
-While in Fashion-MNIST and HAR, the system runs $2,500$ and $1,000$ communication rounds, respectively, and each client only does a single local training iteration per round.
-In our experiments, the root dataset has 100 samples and no intersection with the training and test datasets.
-This setting is similar to FLTrust [9].
-100 is enough for our defense, and it is easy for the server to collect.
-So we do not discuss the influence of the size of the root dataset.
-IV-A6 SkyMask variants
-Intuitively, if the fraction of malicious clients is less than 50% in the real world, we can choose the larger one in two clusters of masks as benign.
-This way, we can further reduce the computational pressure on the server.
-To verify the effectiveness with or without the trusted root model, we consider the following two schemes: SkyMask-NR (no trusted root model): The FL system does not generate a trusted root model, and after all the masks are converged, the system chooses the larger one in two clusters of the masks as benign for aggregation.
-SkyMask: The FL system uses the defense strategy fully following the default steps described in §III-C.
-IV-A7 Metrics
-Our experiments compare the testing accuracy of global models of various FL defenses under different attacks, obtained from the prediction accuracy of the global model on the test dataset.
-The attack success rate under Scaling attack is the fraction of the samples in the test dataset that are identified as the target label after being embedded with triggers.
-To show the malicious client detection ability, we also record false positive rate (FPR) and false negative rate (FNR) in the malicious client detection stage.
-FPR indicates the number of benign clients that are misidentified as malicious clients divided by the total number of benign clients, and FNR indicates the number of malicious clients that are misidentified as benign clients divided by the total number of malicious clients.
-The Mean FPR/FNR is the average of FPR/FNRs in each communication round.
-IV-B The Defense Effectiveness of SkyMask
-We first certificate the normal defensive capability under a low fraction of attacks.
-Then, there is also possibly no attacker in the FL system, so the defense methods should not affect the global model’s performance.
-Eventually, we consider an extreme condition that the fraction of malicious clients is high.
-Figure 4: The impact of high fractions of malicious clients.
-IV-B1 Under a low fraction of attacks
+DBA 0.72/0.99 0.76/0.88 0.76/0.97 0.56/0.09 0.77/0.16 0.77/0.14 0.77/0.11 0.77/0.10
+CIFAR-100 (ResNet20) None 0.44 0.39 0.43 0.17 0.44 0.44 0.44 0.44
+LF 0.41 0.39 0.38 0.11 0.43 0.43 0.44 0.44
+Min-Max 0.16 0.30 0.16 0.05 0.16 0.19 0.44 0.44
+Min-Sum 0.33 0.28 0.33 0.17 0.35 0.34 0.44 0.44
+Fang-Trim 0.01 0.34 0.04 0.15 0.20 0.22 0.44 0.44
+Fang-Krum 0.03 0.37 0.04 0.03 0.03 0.02 0.44 0.44
+Scaling 0.01/1.00 0.36/0.30 0.43/0.98 0.14/0.00 0.44/0.04 0.44/0.08 0.44/0.01 0.44/0.01
+DBA 0.37/0.98 0.40/0.90 0.44/0.94 0.16/0.01 0.45/0.16 0.45/0.15 0.44/0.02 0.44/0.01
+4.2.1 Under a low fraction of attacks.
 By default, we set the fraction of malicious clients to 20% following the default setting in [5].
-Table I shows that, on Fashion-MNIST dataset, Multi-Krum and Bulyan algorithms perform robustness under LF attack, but they fail in the other attacks.
-FLTrust is relatively good that in a few experiments it achieves the same performance as ours in the non-IID case, while under LF attack, it has a 3% accuracy loss.
-Our SkyMask achieves the highest accuracy rate on Fashion-MNIST under different attacks compared to all others.
-Scaling attack does not show an effective impact on all the defense methods.
-To demonstrate the generalization, we conduct the experiments on CIFAR-10 dataset using a larger model, ResNet20.
-In the IID and non-IID cases, FLTrust does not perform as well as it does on FashionMNIST dataset, with a severe drop in accuracy under multiple attacks (especially the Min-Max attack).
-The other defense methods have little or no defensive effect.
-On the contrary, our SkyMask reaches the highest accuracy under attacks.
-In the IID case, our method achieves the most significant gap with other algorithms in Min-Max attack experiments.
-The testing accuracy is up to $0.79$, while the accuracy results obtained by others are less than $0.69$.
-Since the samples with target label 0 are 10% of the test dataset in FashionMNIST and CIFAR-10, the attack success rate is about 10% if there is no attack.
-We can see that FedAvg is overwhelmed and Trim suffers from the Scaling attack severely on CIFAR-10.
-SkyMask still keeps the robustness and performs well.
-The HAR dataset has a real-world data distribution, making the experiment more meaningful and closer to reality.
-We can see in the HAR part of Table I that Trim, Krum, and their variants cannot work well.
-FLTrust always has a little gap with the accuracy of models not under attack, which means it could not eliminate the impact of attacks, but it is still acceptable.
-Unlike all these defense methods, our SkyMask keeps the global model immune to attacks.
-Due to the attack designers’ optimization for the abnormality degree, only some particular parameters of the poisoning model updates changed with emphasis.
-It can be seen that the four fine-grained model poisoning attacks (Min-max, Min-Sum, Fang-Trim, and Fang-Krum) bypass the existing defense methods in many cases.
+Table 1 shows that the testing accuracy of most existing defenses is reduced by more than 10% with at least one attack.
+FLTrust is a relatively strong baseline defense because it achieves similar good performance as ours in a few Fashion-MNIST experiments.
+While under LF attack, FLTrust has a 3% accuracy loss.
+Our SkyMask achieves the highest testing accuracy on Fashion-MNIST under different attacks compared to all others.
+To demonstrate the generalization, we experiment on CIFAR-10 and CIFAR-100 dataset using a larger model, ResNet20.
+In such scenarios, multi-client fine-grained attacks are much more powerful.
+Trim, Krum, DeepSight, and FLAME are out of defense.
+FLTrust does not perform as well as it does on Fashion-MNIST dataset, with a severe drop in accuracy under various attacks.
+In contrast, our SkyMask reaches the highest accuracy under all attacks.
+Especially on CIFAR-100 dataset, our method achieves the most significant gap with other algorithms in Min-Max attack experiments.
+The accuracy is up to $0.44$, while the accuracy results obtained by others are less than $0.3$.
+Besides, SkyMask achieves the best and unattacked-level performance under the targeted attacks.
+Due to the attackers’ optimization for the abnormality degree, only some particular parameters of the poisoning models changed with emphasis.
+The results show that the four fine-grained attacks (Min-max, Min-Sum, Fang-Trim, and Fang-Krum attacks) bypass the existing defense methods in many cases.
 In all these experiments, SkyMask not only achieves the highest testing accuracy but reaches the same accuracy level as unattacked FedAvg’s, i.e., the same accuracy or no more than 1% accuracy loss.
-As Section IV-A6 describes, the server does not need to consume resources to generate a root model if the fraction of malicious clients is less than 50%.
-In the 20% case, as shown in Table I, SkyMask-NR achieves good results under any attack and on any dataset, which shows the same defensive capability as SkyMask.
-Therefore, if the server confirms that the number of malicious clients is less than half the total number, it can choose SkyMask-NR as the defense strategy to reduce the overhead while maintaining the same defensive capability.
-For generalization, SkyMask can be applied to handle more cases.
-IV-B2 Under no attack
-Some attacks send poisoned models to the server only at certain communication rounds, and some model updates are often mistaken for poisoning by existing defense methods due to data heterogeneity.
-The system collapses if the global model performance suffers from the defense strategy, even when there is no attack.
-So when there is no attack, the expected result is that any method should not affect the performance of the basic aggregation algorithm FedAvg can achieve.
-In Table I, almost all other robust aggregation algorithms affect the convergence prediction accuracy in the no-attack case.
-For instance, FLTrust has an accuracy loss of about 2% for the non-IID case on CIFAR-10 dataset, and Krum algorithm has an accuracy loss ranging from about 4% to 20% on different datasets.
-At the same time, our SkyMask maintains a comparable performance as unattacked FedAvg on all datasets with a difference of less than 1%.
+As §4.1.4 describes, the server does not need to consume resources to generate a root model if the fraction of malicious clients is less than 50%.
+In the 20% case, as shown in Table 1, SkyMask-NR achieves good results under any attack and on any dataset, which shows the same defensive capability as SkyMask.
+Therefore, if the server confirms that the number of malicious clients is less than half of the total number, it can choose SkyMask-NR as the defense strategy to reduce the overhead while maintaining the same defensive capability.
+For a more general scenario, it can choose SkyMask as the defense strategy.
+4.2.2 Under no attack.
+Some attacks send poisoned models to the server only at certain rounds, and some model updates can be mistaken for malicious by existing defense methods due to data heterogeneity.
+The system collapses if the global model performance suffers from the defense, even when there is no attack.
+So when there is no attack, the expected result is that any method should not affect the performance achieved by the basic aggregation algorithm FedAvg.
+In Table 1, almost all the existing robust aggregation algorithms affect the testing accuracy in the no-attack case.
+For instance, FLTrust has an accuracy loss of about 2% on CIFAR-10 dataset and 5% on CIFAR-100.
+Krum algorithm has an accuracy loss ranging from about 6% to 27% on different datasets.
+In contrast, our SkyMask maintains a comparable performance as unattacked FedAvg on all datasets with a difference of less than 1%.
 The existing robust aggregation algorithms either choose a subset of model updates that seem benign or use model updates’ statistics to correct the impact.
@@
 If there is no attack, our masks form only one cluster, or there are only several outliers.
-So the server selects most clients and has the same or very similar convergence with FedAvg in both IID and non-IID cases.
-TABLE II: Testing accuracy, FPR and FNR of different malicious client detection methods under different attacks.
-The experimental results of Scaling attack is in the form of “testing accuracy / attack success rate”.
+So SkyMask selects most of clients and has the same or very similar convergence compared with FedAvg.
+Figure 6: The impact of high fractions of malicious clients under fine-grained attacks.
+4.2.3 Under a high fraction of attacks.
+The experiments mentioned in the previous sections set the fraction of malicious clients to 20%, while in a real application environment, an attacker can control a larger fraction of malicious clients for the attack.
+Therefore, to verify the defensive capability of our system in this case, we use four-layer CNN as the global model on Fashion-MNIST dataset and conduct comparison experiments with 40%, 60%, and 80% malicious fractions.
+We present the results pertaining to fine-grained attacks in Fig.
+6.
+Additional results can be found in the supplementary §B.1.
+Only SkyMask and FLTrust maintain the defensive capability under fine-grained attacks.
+The other defense algorithms completely lose their defensive capability, exhibiting a significant gap from the unattacked level.
+From Fig.
+6(b), Fig.
+6(c), and Fig.
+6(d), we can see that SkyMask converges more stable than FLTrust when they are attacked by Fang-Krum, Min-Max, and Min-Sum attacks.
+In Fig.
+6(d), the accuracy obtained by SkyMask is 2% higher than FLTrust’s under Min-Sum attacks.
+When the fraction is more than 50%, only SkyMask maintains the defensive capability.
+Even as the best-performing baseline defense, FLTrust’s performance worsens as the fraction rises.
+FLTrust’s testing accuracy fluctuates more, and the global model even converges in the wrong direction, e.g.
+in Fig.
+6(l), the accuracy obtained by FLTrust decreases from 85% to 81% after a sharp fluctuation.
+Under other attacks, the increasing fraction influences FLTrust’s convergence speed, whereas SkyMask remains unaffected and converges faster than FLTrust.
+Table 2: Testing accuracy, FPR, and FNR of different malicious client detection methods under different attacks.
+The experimental results of targeted attacks are in the form of “testing accuracy/attack success rate.”
 Dataset (Model) Attack Testing accuracy FPR FNR
 Tolpegin FLDetector SkyMask Tolpegin FLDetector SkyMask Tolpegin FLDetector SkyMask
-Fashion -MNIST (CNN) non-IID None 0.89 0.86 0.89 / / / / / /
-LF 0.89 0.86 0.89 0.11% 0.05% 0.00% 0.38% 0.00% 0.02%
-Min-Max 0.88 0.08 0.89 0.38% 96.94% 0.27% 0.80% 100% 0.40%
-Min-Sum 0.64 0.38 0.89 41.1% 100% 0.26% 84.8% 100% 0.00%
-Fang-Trim 0.88 0.86 0.89 0.00% 0.00% 0.00% 0.00% 0.00% 0.00%
-Fang-Krum 0.85 0.10 0.89 28.2% 99.46% 0.20% 60.4% 100% 0.00%
-Scaling 0.89 / 0.10 0.89 / 0.10 0.89 / 0.10 0.00% 0.00% 0.00% 0.03% 0.00% 0.00%
-Fashion -MNIST (CNN) IID None 0.90 0.88 0.90 / / / / / /
-LF 0.89 0.87 0.90 0.07% 0.02% 0.00% 0.00% 0.00% 0.00%
-Min-Max 0.89 0.87 0.90 0.06% 0.00% 0.00% 0.00% 0.81% 0.00%
-Min-Sum 0.88 0.49 0.90 15.34% 100% 0.01% 7.60% 100% 0.00%
-Fang-Trim 0.90 0.88 0.90 0.00% 0.00% 0.00% 0.00% 0.00% 0.00%
-Fang-Krum 0.89 0.13 0.90 8.04% 99.60% 0.19% 6.40% 100% 0.00%
-Scaling 0.89 / 0.11 0.90 / 0.11 0.90 / 0.10 0.00% 0.00% 0.00% 0.68% 0.00% 0.00%
-CIFAR10 (ResNet20) non-IID None 0.76 0.75 0.76 / / / / / /
+CIFAR10 (ResNet20) None 0.76 0.75 0.76 0.00% 18.2% 0.00% / / /
 LF 0.70 0.72 0.76 13.1% 0.03% 4.72% 19.9% 25.53% 2.60%
@@
 Scaling 0.76/0.10 0.74/0.10 0.77/0.11 0.00% 0.00% 0.00% 0.10% 0.00% 0.00%
-CIFAR10 (ResNet20) IID None 0.75 0.79 0.79 / / / / / /
-LF 0.73 0.75 0.78 3.45% 0.00% 3.24% 5.80% 4.26% 0.00%
-Min-Max 0.74 0.25 0.79 0.05% 100% 0.00% 0.00% 100% 0.00%
-Min-Sum 0.74 0.31 0.79 0.14% 100% 0.00% 1.00% 100% 0.00%
-Fang-Trim 0.74 0.77 0.78 0.03% 0.00% 0.00% 0.00% 0.05% 0.00%
-Fang-Krum 0.74 0.18 0.79 4.04% 94.60% 0.00% 10.60% 100% 0.00%
-Scaling 0.79 / 0.09 0.78 / 0.10 0.79 / 0.10 0.00% 0.00% 0.00% 0.00% 0.00% 0.00%
-HAR (LR) None 0.96 0.92 0.96 / / / / / /
-LF 0.96 0.68 0.96 0.22% 13.27% 0.00% 0.13% 100% 0.00%
-Min-Max 0.96 0.07 0.96 1.38% 100% 0.83% 2.00% 100% 0.00%
-Min-Sum 0.95 0.53 0.96 31.6% 100% 4.53% 32.0% 100% 4.00%
-Fang-Trim 0.95 0.96 0.96 0.00% 0.00% 0.96% 0.00% 0.00% 0.00%
-Fang-Krum 0.95 0.13 0.96 26.4% 100% 0.03% 29.3% 100% 0.00%
-Scaling 0.96 / 0.20 0.95 / 0.19 0.96 / 0.20 0.25% 0.00% 0.00% 12.58% 0.00% 0.08%
-IV-B3 Under a high fraction of attacks
-The experiments mentioned in the previous sections set the fraction of malicious clients to 20%, while in a real application environment, an attacker can control a larger fraction of malicious clients for the attack.
-Therefore, to verify the defensive capability of our system in this case, we use CNN as the global model on Fashion-MNIST dataset in the non-IID case and conduct comparison experiments with 40%, 60%, and 80% of malicious clients.
-Since Bulyan requires that the number of malicious clients $m$ should satisfy $n\geq 4m+3$, these experiments do not include this algorithm.
-Trim requires that the number of malicious clients should meet $n\geq 2m+1$, and Multi-Krum requires that $n-c>2m+2$, so these aggregation algorithms are only included in 40% experiments.
-The experimental results are shown in Figure 4.
-Most algorithms perform well under Scaling attack except for FedAvg and Krum.
-When the fraction of malicious clients is 40%, Trim loses its defensive capability completely, and Multi-Krum can only work under LF attack.
-Krum can reduce the impact of attacks under LF and Fang-Trim attacks, but the accuracy loss is still large.
-Although the accuracy of FLTrust fluctuates in some communication rounds, it still tends to converge to an unattacked level.
-From Figure 4(c), Figure 4(d) and Figure 4(e), we can see that SkyMask converges more stable than FLTrust when they are attacked by Fang-Krum attack, Min-Max attack and Min-Sum attack.
-In Figure 4(a) and Figure 4(e), the accuracy obtained by SkyMask is 2% to 4% higher than FLTrust’s under LF and Min-Sum attacks.
-When the fraction is more than 50%, only FLTrust and SkyMask maintain the defensive capability, but FLTrust’s performance worsens as the fraction rises.
-FLTrust’s testing accuracy fluctuates even more, and the global model even converges in the wrong direction (e.g., in Figure 4(q), the accuracy obtained by FLTrust decreases from 85% to 81% after a sharp fluctuation).
-Under other attacks, the increasing fraction influences FLTrust’s convergence speed, and SkyMask converges faster than it.
-(a) Fang-Krum attack, Fashion-MNIST
-(b) Min-Sum attack, Fashion-MNIST
-(c) LF attack, CIFAR-10
-Figure 5: The impact of detection failure and the training efficiency of Tolpegin defense, FLDetector and SkyMask.
-(a) On FashionMNIST Dataset (non-IID) with CNN
-(b) On FashionMNIST Dataset (IID) with CNN
-(c) On CIFAR-10 Dataset (non-IID) with ResNet20
-(d) On CIFAR-10 Dataset (IID) with ResNet20
-Figure 6: The impact of root data distribution on SkyMask.
-IV-C The Significance of Learnable Masks
-We conduct comparative experiments on SkyMask and the other malicious client detection algorithms, i.e., Tolpegin defense [2] and FLDetector [16] (combined with Median).
-We sample the result of malicious client detection every ten communication rounds and calculate all the metrics described in §IV-A7.
-In Table II, we can see that FLDetector only works under LF attack, Fang-Trim attack and Scaling attack.
+DBA 0.65/0.47 0.77/0.10 0.77/0.10 0.00% 0.00% 0.00% 46.1% 0.00% 0.00%
+4.3 The Significance of Learnable Masks
+We conduct comparative experiments on SkyMask and the other malicious client detection algorithms, i.e., Tolpegin defense [30] and FLDetector [36].
+We sample the result of malicious client detection every ten communication rounds and calculate all the metrics described in §4.1.5.
+In Table 2, we can see that FLDetector only works under LF attack, Fang-Trim attack and targeted attack.
 Min-Max and Min-Sum attacks make FLDetector completely confuse the malicious model updates with the benign model updates.
-In the Fashion-MNIST part of Table II, we can see that it is hard for Tolpegin defense to detect malicious clients stably, especially under Fang-Krum attack and Min-Sum attack.
-Its FPR and FNR are pretty high, so it fails to defend against attacks effectively and loses a lot of benign information.
-Quite the opposite, SkyMask perfectly completes the detection task and achieves an FNR of less than 4‰  on Fashion-MNIST dataset with a CNN model, which means that the server hardly misses any malicious client, and it can eliminate the harm of attacks.
-SkyMask also achieves an FPR of less than 3‰ , so it can preserve most benign clients and better characterize the global data distribution.
-Figure 5(a) and Figure 5(b) show the trend of testing accuracy on non-IID Fashion-MNIST with a CNN global model under Fang-Krum attack and Min-Sum attack.
-Detection failure means that the FL defense system does not detect all the malicious clients in that round.
-We can see that SkyMask gradually rises and converges without detection failure.
-On the contrary, Tolpegin defense converges but with poor testing accuracy under Fang-Krum attack, and it is damaged under Min-Sum attack.
-Each detection failure causes a fluctuation in testing accuracy.
-In the CIFAR-10 part of Table II, Tolpegin defense performs even worse in the non-IID case with large model architecture (ResNet20).
-Except for Fang-Trim attack, all other attacks greatly harm it.
+Tolpegin defense only survives under Fang-Trim attack, and all other attacks greatly harm it.
 The large FNR of Tolpegin defense means many malicious clients complete their attacks.
 The large FPR of Tolpegin defense shows that the remaining benign clients may lose some representation, so Tolpegin defense achieves a lousy performance.
-In the IID case, Tolpegin defense resumes partial defensive capability, but there is still about 10% FNR under Fang-Krum attack.
-However, our SkyMask achieves the same prediction accuracy as a no-attack case, with FPR less than 5% and FNR less than 3%.
-LF attack on CIFAR-10 dataset is more stealthy, and the FNR of SkyMask rises to 2.6%.
-As Figure 5(c) shows, at the beginning of FL training under LF attack, determining whether the difference between clients comes from data distribution or malicious client’s anomaly operation is more challenging.
-However, this does not matter since the LF attack is weak at the beginning of the training process.
-The global model still moves towards convergence; after several communication rounds, SkyMask can resume normal detection ability and identify these malicious clients.
-This mistake does not prevent our method from helping the global model reach convergence and high testing accuracy.
-TABLE III: The impact of threshold in binarization on SkyMask.
-Dataset (Model) Attack Acc. FPR FNR
-0.3 0.4 0.5 0.6 0.7 0.3 0.4 0.5 0.6 0.7 0.3 0.4 0.5 0.6 0.7
-Fashion -MNIST (CNN) non-IID None 0.89 0.89 0.89 0.89 0.89 / / / / / / / / / /
-LF 0.88 0.88 0.89 0.89 0.88 0.01% 0.02% 0.00% 0.03% 0.32% 0.08% 0.02% 0.02% 0.00% 0.04%
-Min-Max 0.89 0.89 0.89 0.89 0.89 0.00% 0.00% 0.27% 0.00% 0.00% 0.00% 0.00% 0.40% 0.00% 0.00%
-Min-Sum 0.88 0.89 0.89 0.89 0.88 0.04% 0.20% 0.26% 0.34% 0.89% 0.00% 0.40% 0.00% 0.00% 1.60%
-Fang-Trim 0.88 0.89 0.89 0.88 0.88 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00%
-Fang-Krum 0.89 0.89 0.89 0.89 0.89 0.18% 0.28% 0.20% 0.40% 0.28% 0.40% 0.00% 0.00% 0.40% 0.00%
-Scaling 0.89 / 0.10 0.89 / 0.12 0.89 / 0.10 0.89 / 0.10 0.89 / 0.11 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00%
-Fashion -MNIST (CNN) IID None 0.88 0.89 0.90 0.89 0.89 / / / / / / / / / /
-LF 0.89 0.89 0.90 0.90 0.89 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00%
-Min-Max 0.89 0.90 0.90 0.90 0.89 0.08% 0.00% 0.00% 0.00% 0.01% 0.00% 0.00% 0.00% 0.00% 0.00%
-Min-Sum 0.89 0.89 0.90 0.90 0.89 0.04% 0.05% 0.01% 0.01% 0.34% 0.13% 0.00% 0.00% 0.00% 1.00%
-Fang-Trim 0.89 0.90 0.90 0.90 0.90 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00%
-Fang-Krum 0.89 0.89 0.90 0.90 0.89 0.77% 0.56% 0.19% 0.05% 0.07% 1.00% 0.20% 0.00% 0.00% 0.20%
-Scaling 0.89 / 0.12 0.90 / 0.11 0.90 / 0.10 0.90 / 0.11 0.90 / 0.10 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00%
-CIFAR-10 (ResNet20) non-IID None 0.76 0.76 0.76 0.76 0.76 / / / / / / / / / /
-LF 0.76 0.76 0.76 0.76 0.76 4.72% 4.33% 2.03% 2.40% 6.65% 5.70% 2.80% 2.60% 4.00% 6.30%
-Min-Max 0.77 0.77 0.77 0.77 0.77 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00%
-Min-Sum 0.77 0.77 0.77 0.77 0.77 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00%
-Fang-Trim 0.76 0.77 0.76 0.76 0.76 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00%
-Fang-Krum 0.77 0.77 0.77 0.77 0.77 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00%
-Scaling 0.76 / 0.09 0.76 / 0.12 0.77 / 0.12 0.77 / 0.10 0.77 / 0.10 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00%
-CIFAR-10 (ResNet20) IID None 0.79 0.78 0.79 0.78 0.78 / / / / / / / / / /
-LF 0.76 0.78 0.78 0.79 0.78 5.31% 3.40% 3.24% 1.25% 8.04% 5.45% 1.75% 0.00% 4.45% 8.45%
-Min-Max 0.78 0.79 0.79 0.79 0.78 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00%
-Min-Sum 0.78 0.79 0.79 0.79 0.79 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00%
-Fang-Trim 0.78 0.79 0.78 0.78 0.78 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00%
-Fang-Krum 0.78 0.79 0.79 0.78 0.78 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00%
-Scaling 0.79 / 0.11 0.79 / 0.10 0.79 / 0.10 0.79 / 0.09 0.79 / 0.11 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00%
-HAR (LR) None 0.96 0.96 0.96 0.96 0.96 / / / / / / / / / /
-LF 0.96 0.96 0.96 0.96 0.96 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00%
-Min-Max 0.96 0.96 0.96 0.96 0.96 1.70% 1.57% 0.97% 2.50% 0.43% 1.00% 0.00% 1.00% 1.00% 1.40%
-Min-Sum 0.96 0.96 0.96 0.96 0.96 2.17% 2.63% 0.87% 3.60% 5.70% 2.50% 2.00% 0.70% 1.00% 3.00%
-Fang-Trim 0.96 0.96 0.96 0.96 0.96 0.16% 0.00% 0.00% 0.00% 0.00% 2.63% 0.00% 0.00% 0.00% 0.00%
-Fang-Krum 0.96 0.96 0.96 0.96 0.96 0.70% 0.60% 0.00% 0.07% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00%
-Scaling 0.96 / 0.22 0.96 / 0.19 0.96 / 0.20 0.96 / 0.16 0.96 / 0.20 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00%
-Figure 7: The impact of root data distribution on model convergence.
-For HAR dataset, Tolpegin defense’s testing accuracy seems good even when FNR rises to about 30%.
-The reason is that since the classification task is simple, even if detection failure in some rounds leads to the fluctuation of the testing accuracy, the testing accuracy can still rise to a reasonable level through successive rounds of detection success.
-In contrast, SkyMask achieves a good malicious client detection result with FNR less than 4%, maintaining high testing accuracy.
-IV-D Sensitivity analysis
-(a) On FashionMNIST Dataset (non-IID) with CNN
-(b) On FashionMNIST Dataset (IID) with CNN
-(c) On CIFAR-10 Dataset (non-IID) with ResNet20
-(d) On CIFAR-10 Dataset (IID) with ResNet20
-(e) On HAR Dataset with LR
-Figure 8: The impact of dimensionality reduction on SkyMask.
-IV-D1 The Impact of Root Dataset Distribution
-The server’s root dataset is randomly sampled from the global data distribution by default.
-It is most likely to collect a root dataset with bias in the real world.
-However, we can prove that as soon as the bias is not too large ($p\leq 0.8$), our method can still take advantage of the root dataset and complete the malicious client detection task.
-We experiment on Fashion-MNIST with a CNN global model and CIFAR-10 with a Resnet20 model.
-The results are as testing accuracy part of Figure 6 shown: most of the testing results of the global model obtained by our method are very close or equal to the model obtained by FedAvg under no attack.
-However, in some cases, such as Figure 6(b) under Min-Max attack, and Figure 6(d) under no attack and LF attack, the testing accuracy is not acceptable when the bias probability of the root dataset is 0.95.
-Considering that the root dataset has a total of 100 samples, the lack of some classes of samples in the root dataset is possible when the bias probability of the root dataset is 0.95, so this result is reasonable.
-The FPR and FNR results in the figures show that the malicious client detection accuracy of SkyMask fluctuates as the degree of non-IID of the root data changes but remains very low.
-Our method does not fail until the bias probability of the root dataset exceeds 80%.
-As long as the generated root dataset has a bias not so significant, SkyMask can complete its malicious client detection task and keep the global model a good performance.
-In Figure 7, we show the variation curve of accuracy of SkyMask for non-IID FashionMNIST experiments.
-The convergence speed reduces along with the data distribution bias increasing, especially under Fang-Trim attack and Scaling attack.
-Although the root dataset’s different bias probability ($p\leq 0.8$) does not affect the convergence result, this phenomenon tells us that the imbalanced root dataset can slow down the global model’s convergence, and we had better choose a good root dataset for the server.
-IV-D2 The Impact of Thresholds in Binarization
-As the description of the processing of binary masks in Section III-C, the server should choose a threshold $\tau$.
-Our optimization strategy makes the parameters in masks converge to 0 or 1, so SkyMask should not be sensitive to threshold $\tau$.
-We conduct experiments with different thresholds from 0.3 to 0.7, and the results are shown in Table III.
-On Fashion-MNIST dataset, it can be seen that SkyMask loses a little accuracy under some attacks when the threshold goes far away from 0.5.
-Still, it also has a similar accuracy as unattacked FedAvg’s level, with about 1% accuracy loss.
-SkyMask achieves FNR lower than 1.6%, so there are almost no undetected malicious clients; FPR is lower than 0.9%, so most benign clients are retained.
-On CIFAR-10 dataset, especially under LF attack, FPR and FNR increase as the threshold deviates from 0.5.
-But they are still less than 9%, and the testing accuracy loss is less than 1%.
-And for HAR, we can see similar results that SkyMask keeps a very low FPR and FNR, achieving the same testing accuracy as unattacked FedAvg.
-Therefore, the change in threshold does not affect our system’s performance, and SkyMask is not sensitive to the threshold in binarization.
-IV-D3 The Impact of Dimension Reduction
-The clustering results of SkyMask are influenced by the dimensionality of the features after dimensionality reduction.
-In Figure 8, we can see that when the number of features after dimensionality reduction rises from 2 to 15, although the testing accuracy does not decrease severely in most cases, there is a noticeable increase in the FPR and FNR, especially under LF attack.
-The reason is that the framework uses the Gaussian mixture model to cluster the dimension-reduced masks.
-Since the total number of masks is not large enough, the clustering results may not be satisfying in high-dimension space due to the curse of dimensionality.
-Therefore, we set the number of features after dimensionality reduction as two empirically.
-IV-E SkyMask’s Scalability
-To demonstrate the scalability of SkyMask, we conducted experiments utilizing a CNN and ResNet20 global model trained on Fashion-MNIST and CIFAR-10, respectively.
-For Fashion-MNIST, we assessed the prediction accuracy of the main task, along with the FPR and FNR of SkyMask under various attack scenarios, encompassing 500 and 1000 clients.
-We have mentioned in §IV-A1 that the local datasets assigned to the clients were disjoint and each client’s dataset should be reasonably sized, so for the CIFAR-10 experiments with fewer data samples, we selected 200 and 500 clients.
-Table IV shows that SkyMask can work well even if the number of clients increases to 1000 on Fashion-MNIST.
-As the number of clients increases, more malicious clients are missed under Min-Sum and Fang-Krum attack, but the FNR is still lower than 1.6%.
-In addition, the missed malicious local models do not affect the performance of the global model.
-We can see that the global model of SkyMask achieves the same testing accuracy as the unattacked global model’s obtained by FedAvg.
-For the experiments on CIFAR-10, we observed a decrease in the testing accuracy results of the unattacked FedAvg as the total number of clients increased.
-This can be attributed to the challenges posed by the smaller size of the local datasets and the larger number of clients in the FL training process [23].
-However, our SkyMask algorithm demonstrated robust performance even under most attacks.
-For the LF attack, SkyMask achieved comparable testing accuracy to the unattacked FedAvg, despite a slight increase in the FNR to approximately 9%.
-Therefore, SkyMask can work well when the total number of clients increases.
-Even if there are millions of clients in the FL system, the server can randomly choose a fixed number (e.g., 100 or 500) of clients in each communication round to execute malicious client detection and aggregation.
-The details we will not describe in this paper.
-TABLE IV: The impact of the total number of clients. (The first line in table is the testing accuracy result of FedAvg under no attack.)
-FashionMNIST (CNN) Test Acc. FPR FNR
-500 1000 500 1000 500 1000
-FedAvg 0.89 0.89 / / / /
-None 0.89 0.89 / / / /
-LF 0.89 0.89 0.40% 0.40% 0.94% 0.62%
-Min-Max 0.89 0.89 0.40% 0.30% 0.00% 0.00%
-Min-Sum 0.89 0.89 0.40% 1.60% 0.00% 0.00%
-Fang-Trim 0.89 0.89 0.00% 0.00% 0.00% 0.00%
-Fang-Krum 0.89 0.89 0.60% 1.20% 0.00% 0.00%
-Scaling 0.89 / 0.11 0.89 / 0.10 0.00% 0.00% 0.00% 0.00%
+Our SkyMask achieves the same prediction accuracy as a no-attack case, with FPR less than 5% and FNR less than 3%.
+Fig.
+LABEL:fig:detect_a and Fig.
+LABEL:fig:detect_b show the trend of testing accuracy on Fashion-MNIST with a four-layer CNN global model under Fang-Krum and Min-Sum attack.
+Detection failure means that the FL defense system does not detect all the malicious clients in that round.
+Under Fang-Krum attack, Tolpegin defense converges but yields poor testing accuracy, while FLDetector is rendered ineffective.
+Additionally, both FLDetector and Tolpegin defense exhibit fluctuating and low testing accuracy under Min-Sum attack due to detection failures.
+In contrast, SkyMask stably converges without encountering any detection failures.
+Table 3: The impact of the total number of clients. (The first line in table is the testing accuracy result of FedAvg under no attack.)
 CIFAR-10 (ResNet20) Test Acc. FPR FNR
@@
 Scaling 0.75/0.09 0.74/0.11 0.00% 0.00% 0.00% 0.00%
-V Discussion and Limitations
-The current Byzantine attacks are launched with either a single client or multiple clients executing the same attack strategy, which may only account for an increased proportion of the model aggregation stage on the server, but the malicious clients do not collaborate.
-Increasingly, some researchers find that some specific layers are more vulnerable to Byzantine attack [24].
-We also observe that some Byzantine attacks modify parameters near the output layer significantly more than those far from the output layer.
-Therefore, each client can be responsible for a different part of model poisoning, using the philosophy of “not putting eggs in the same basket.” The attacker hides poisoning factors in all the uploaded models of the malicious clients it compromises.
-When most or all malicious model updates bypass the malicious client detection mechanism and get aggregated on the server, it can form a powerful attack on the global model, which is truly a collaborative attack by multiple malicious clients.
-The damage to the characteristics of local model updates by this collaborative attack is unknown, and it remains to be verified whether SkyMask can efficiently detect these malicious attacks.
-For other malicious client detection algorithms, detecting this abnormal pattern in a single model update may be challenging.
-Theoretically, using our mask mechanism, the collaborative impact of all malicious clients on the global model could be shown after masked aggregation so that the optimization process can change their corresponding masks’ pattern.
-Actually, Fang attack and AGR-agnostic attack have attempted to attack collaboratively, which weakens the anomaly level of model updates uploaded by each malicious client but influences the performance of the global model when all attacks are combined.
-Hence, there is a chance for our SkyMask to detect this kind of malicious client.
-In the future, we will continue to study this topic in-depth, first verify the effectiveness of this collaborative attack, and deepen the ability of our learnable binary mask to extract the features of model updates and realize the defense capability for this unknown attack.
-In addition to collaborative attack, the performance of SkyMask under multiple attack types in one communication round is unknown.
-We should explore how many clusters will the corresponding masks obtained by SkyMask form, and we should also adapt the settings of SkyMask according to this.
-Furthermore, SkyMask can be seen as an additional module to the aggregation algorithm.
-Not only the basic FedAvg algorithm, the SOTA FL algorithms, such as FedProx [25], FedBN [26], or FedAMP [27], can also be equipped with our malicious client detection module.
-This feature lets the FL system retain the heterogeneity-handling, high convergence rate, or personalization features the original aggregation algorithm may have.
-It is also possible to enjoy the defensive capability of our malicious client detection strategy, and we will explore the merits of various combinations in future work.
-VI Conclusion
+DBA 0.75/0.10 0.73 /0.10 0.00% 0.00% 0.00% 0.00%
+4.4 SkyMask’s Scalability
+To demonstrate the scalability of SkyMask, we conduct experiments utilizing a ResNet20 global model trained on CIFAR-10.
+We assess the prediction accuracy of the main task, along with the FPR and FNR of SkyMask under various attack scenarios, encompassing 200 and 500 clients.
+In Table 3, we observe a decrease in the testing accuracy results of the unattacked FedAvg as the total number of clients increased.
+This can be attributed to the challenges posed by the smaller size of the local datasets and the larger number of clients in the FL training process [12].
+For the LF attack, SkyMask achieves comparable testing accuracy to the unattacked FedAvg, despite a slight increase in the FNR to approximately 9%.
+For all fine-grained attacks and targeted attacks, SkyMask maintains FPR and FNR at 0%.
+Our SkyMask algorithm demonstrates robust performance under all attacks even when the total number of clients significantly increases.
+5 Conclusion
 We propose a new attack-agnostic robust FL framework called SkyMask to defend against Byzantine attacks.
-By training parameter-level learnable binary masks on a clean root dataset, SkyMask is the first to implement a fine-grained detection of the poisoned elements of local model updates.
-Because SkyMask analyzes the mask as a whole, it avoids accidentally removing outliers caused by data heterogeneity on some elements of model updates and precisely removes the malicious clients.
-Extensive experiments on IID and non-IID data under different attacks prove the effectiveness of SkyMask.
-Our method shows a solid defensive capability, better than various robust aggregation algorithms and existing malicious client detection methods.
-Moreover, it also tackles the problem that most clients are malicious and can defend against attacks with a fraction of malicious clients up to $80\%$.
+By training parameter-level learnable binary masks on a clean root dataset, SkyMask is the first to design a fine-grained detection of the poisoned elements of local model updates.
+Extensive experiments on non-IID datasets under different attacks prove the effectiveness, generality, and scalability of SkyMask.
+Our SkyMask shows a solid defensive capability, better than various robust aggregation algorithms and existing malicious client detection methods.
+Moreover, our SkyMask also tackles the problem that most clients are malicious and can defend against attacks with a fraction of malicious clients up to $80\%$.
+Acknowledgements
+This work was partially supported by the National Key R&D Program of China (2022YFB4402102), the Shanghai Key Laboratory of Scalable Computing and Systems, the HighTech Support Program from STCSM (No.22511106200), and Intel Corporation (UFunding 12679).
+Tao Song is the corresponding author.